### v4

# A Few Large Shifts: Layer-Inconsistency Based Minimal Overhead Adversarial Example Detection

Abstract
[…]
Keywords Adversarial detection  $\cdot$ Deep learning  $\cdot$ Layer inconsistency  $\cdot$ Robust defense  $\cdot$ Adaptive attacks
1 Introduction
Figure 1: Illustration of A Few Large Shifts Assumption–an assumption on a few large perturbation shifts on a simple target classifier causes vulnerable manifolds that can trigger an unexpected transition to a different manifold space. Leveraging these vulnerable regions, an adversarial sample is crafted using an attack algorithm that applies perturbation noise, causing the original sample to shift from the cat class manifold to the dog class manifold through transitions occurring at specific targeted layers.
Deep neural networks (DNNs) have been broadly deployed in computer vision, natural language processing, multi-modal tasks, and beyond [26, 16, 37]. Although they exhibit remarkable performance, adversarial examples (AEs)—inputs containing small yet malicious perturbations—can induce misclassifications while appearing virtually unchanged to human observers [8]. This vulnerability poses severe risks in high-stakes domains such as autonomous driving or disease diagnosis [38, 28], underlining the need for robust defenses.
Existing defensive strategies generally fall into three categories: adversarial training, input purification, and AE detection [13, 1]. Adversarial training, re-trains or fine-tunes a model using adversarially perturbed samples [34, 20], despite its robustness [10, 36], it incurs high computational costs and may compromise clean-data accuracy. Input purification methods [24, 30] attempt to remove adversarial noise through preprocessing (e.g., denoising), yet often fail against adaptive attacks [7].
A more flexible alternative is AE detection, which avoids retraining or significantly modifying the model by rejecting suspicious inputs at test time [35, 40]. Adversarial detection methods can be broadly categorized into two types: those that compare inputs to a reference set or distribution of known patterns (e.g., Deep k-Nearest Neighbors, DkNN [27]; Latent Neighborhood Graph, LNG [1]), and those that analyze invariants in the learned representations [15, 6]. Reference-based detectors often require storing adversarial examples or constructing complex neighbor graphs, which can be computationally intensive and prone to overfitting specific attack types. To address these limitations, recent work [39] proposed detecting consistency between augmented inputs using pre-trained Self-Supervised Learning (SSL) models. Despite avoiding both the need for complex data structures and reliance on adversarial examples, it introduces several practical challenges. Notably, it incurs significant overhead due to the use of large pre-trained SSL models–often significantly larger (e.g., $\times2$ larger params) than the target classifier itself–and assumes the availability of high-quality SSL models for each domain, which may not always be feasible. Furthermore, their method relies solely on SSL models pre-trained on the same datasets used for evaluation.
To eliminate the overhead of external models, complex data structures, or heavy augmentations, we develop a self-contained, layer-wise detector that scrutinizes the network’s own representations. This raises a central question: How do adversarial perturbations propagate through a deep neural network? Prior work shows that such perturbations often leave early-layer features largely intact while inducing sharp deviations in deeper layers [15]. Rather than contrasting these deviations with external reference sets or heavy augmented features, we ask our core research question: Do manifold-induced feature jumps themselves expose adversarial inputs? We formalize this intuition in the A Few Large Shifts Assumption, which posits that adversarial perturbations push representations onto fragile submanifolds at only a few critical layers, producing large, localized shifts. As illustrated in Figure 1, these sparse disruptions leave a detectable footprint across successive layers. Guided by this assumption, we introduce a lightweight detection framework grounded in layer-wise inconsistency. Two complementary probes—Recovery Testing (RT) for intermediate activations and Logit-layer Testing (LT) for the output layer—are calibrated solely on benign data. We then fuse their signals through Recovery-and-Logit Testing (RLT), aligning RT and LT scores via quantile normalization to robustly flag a broad spectrum of attacks, including those that evade one probe in isolation.
The distinguishing features and requirements of our method are summarized in Table 1, where we compare RT and LT against several representative baselines. As shown, our method is self-sufficient (requires no adversarial examples), model-local (requires no external SSL), and low-overhead (no kNN graphs or excessive augmentation), making it suitable for scalable deployment.
Our key contributions are as follows:
  • We propose a novel layer-wise adversarial detection paradigm that exploits partial consistency in internal representations of adversarial examples by focusing on feature transformations. Unlike prior methods, our approach avoids significant overhead by eliminating the need for heavy augmentations, complex structures to maintain reference sets, external large pre-trained models, such as SSL, or large ensemble models.
  • We introduce two practical testing strategies—Recovery Testing and Logit-layer Testing—which identify inconsistencies in intermediate and final-layer representations, respectively, using only the target model’s own feature space.
  • We evaluate our framework on standard benchmarks (CIFAR-10, CIFAR-100, and ImageNet), demonstrating strong detection performance, low computational overhead, and negligible impact on clean inputs, making it practical for real-world deployment.
By focusing on unique internal feature transformations over reference-based comparisons or heavy augmentations, our method avoids extensive retraining and specialized pipelines while delivering a more efficient, robust detector. Next, we review related work, detail our layer-consistency analysis, and present experiments across diverse adversarial threat models.
Table 1: Comparison of Requirements / Properties Across Methods. RT = Recovery Testing, LT = Logit-layer Testing. We also include BEYOND [39], LID [23], Mao [24], Hu [14], DkNN [27], kNN-Def. [9]. A   indicates that the method satisfies the criterion, while a   indicates it does not.
Ours Baseline Methods Criterion RT LT BEYOND [39] LID [23] Mao [24] Hu [14] DkNN [27] kNN-Def. [9] Self-Sufficient Data? (No Adversarial Data Needed?) Standalone Model? (No Extra Pre-trained Model Dependencies?) No Heavy kNN Retrieval? (No Nearest-Neighbor Search Overhead?) No Excessive Augmentations? (No Many Augmentations?) No Extra Optimization? (No Additional Training Needed?) /
2 Related Work and Background
Input Transformation-Based Detectors.
A well-established line of work detects adversarial examples by applying simple input transformations and monitoring the model’s response. For instance, [35] proposed feature squeezing, which reduces input precision or applies spatial smoothing to suppress adversarial noise. [18] used quantization and denoising techniques, while other works explore random crops or region replacements. These preprocessing-based detectors are lightweight and model-agnostic but assuming that perturbations are fragile to such transformations. Their effectiveness often hinges on tuning a detection threshold, and their robustness degrades against adaptive attacks that are explicitly trained to remain invariant under the applied transformations. Coarse transformations may also degrade clean-data performance, increasing false positives.
Feature Statistics-Based Detectors.
Another category of detectors analyzes statistical anomalies in the network’s hidden representations. [23] proposed Local Intrinsic Dimensionality (LID) to measure how local feature neighborhoods expand under adversarial perturbations. [11] leveraged kernel density estimation and uncertainty measures, while [17] introduced a Mahalanobis distance-based detector by modeling class-conditional distributions in intermediate layers. These approaches rely on hand-crafted features from latent activations and often require storing high-dimensional embeddings or computing pairwise distances, which can be computationally intensive. Moreover, many methods are vulnerable to adaptive attacks that mimic the distributional properties of benign inputs.
Model Behavior-Based Detectors.
Some methods probe the internal behavior of the model itself to identify inconsistencies caused by adversarial perturbations. [21] encoded final-layer activations into binary vectors for SVM-based detection, while [25] appended auxiliary classifiers to intermediate layers. Others like [22] and [5] focused on monitoring neuron activation paths or ensemble agreement. These methods often require additional training, architectural modifications, or model ensembles, increasing overhead and complicating deployment. Many also depend on adversarial examples for supervision, which limits their generalizability to unseen attacks.
Self-Supervised and Consistency-Based Detectors.
Recent approaches have explored feature consistency using self-supervised learning (SSL) models. For example, BEYOND [39] detects adversarial inputs by measuring feature stability across augmentations in a large SSL encoder. These methods eliminate the need for adversarial examples and external graph structures but introduce high computational costs due to reliance on large pretrained models, which may be significantly larger than the target classifier itself (e.g., $\times2$ larger parameters). Additionally, they often require access to domain-specific SSL models and retraining or fine-tuning of auxiliary heads.
Our Approach.
In contrast to the methods above, our framework proposes a lightweight, attack-agnostic detection mechanism that operates directly on the internal representations of a pre-trained target classifier. It builds on the assumption that adversarial perturbations typically cause large deviations at a few critical layers (A Few Large Shifts Assumption), Our method eliminates the need for adversarial examples, auxiliary models, and graph-based structures by relying solely on benign calibration data and a handful of lightweight augmentations (< 6). By emphasizing feature evolution across layers rather than comparing to external references, we realize a simple, highly efficient detector. Extensive evaluations on CIFAR-10, CIFAR-100, and ImageNet demonstrate strong detection performance with efficient and robust defense against both adaptive and white-box attacks.
Figure 2: Overview of our layer-wise adversarial detection framework. (Left) Recovery Testing (RT) trains regressors $R^{(L\rightarrow k)}$ to reconstruct intermediate features $z_{k}$ from the final embedding $z_{L}$; detection is based on the entropy-weighted reconstruction error $\|z_{k}-R^{(L\rightarrow k)}(z_{L})\|_{2}^{2}$. (Right) Logit-layer Testing (LT) applies perturbations $W^{(g)}$ to compute feature and logit discrepancies, $\Delta z^{(g)}=\|z_{i}-z_{i}^{(g)}\|_{2}^{2}$ and $\Delta\ell^{(g)}=\|o(\sigma(\ell))-\sigma(\ell^{(g)})\|_{2}^{2}$, which are combined as $\frac{H(\sigma(\ell))\Delta\ell^{(g)}}{\Delta z^{(g)}}$. Final RT and LT scores are averaged over $g$. (Bottom) Recovery and Logit Testing (RLT) integrates RT and LT scores using a linear combination after quantile normalization.
3 Methodology
Algorithm 1 Layer-wise Adversarial Detection Measures via RT, LT, and RLT
$f=f_{logit}\circ f_{L}\circ\dots\circ f_{1}$: Target network
2:$\{R^{(L\rightarrow k)}\}_{k=k_{RT}}^{L-1}$: Trained inverse regressors
3:$\{W^{(g)}\}_{g=1}^{G}$: Learned augmentation matrices
4:^$\hat{\mathcal{F}}_{RT},\hat{\mathcal{F}}_{LT}$: Empirical CDFs of RT and LT (from benign data)
5:$\Phi^{-1}$: Standard normal quantile function
6:$x$: Test input
7:
8:function RT($x$)
9:     for $k=k_{RT}$ to $L-1$ do
10:         $e_{k}\leftarrow\|z_{k}(x)-R^{(L\rightarrow k)}(z_{L}(x))\|_{2}^{2}$
11:     end for
12:     $\bm{e}\leftarrow(e_{1},\dots,e_{L-1})$
13:     return $(\log(L-1)-\mathcal{H}(\sigma(\bm{e})))\cdot\log\left(\frac{1}{L-1}\sum_{k}e_{k}\right)$
14:end function
15:
16:function LT($x$)
17:     for $g=1$ to $G$ do
18:         $\Delta z^{(g)}\leftarrow\frac{1}{L-k_{LT}+1}\sum_{i=k_{LT}}^{L}\|z_{i}(x)-z_{i}(W^{(g)}x)\|_{2}^{2}$
19:         $\hat{y}\leftarrow\arg\max\sigma(f(x))$
20:         $\Delta\ell^{(g)}\leftarrow\|\mathbf{o}_{\hat{y}}-\sigma(f(W^{(g)}x))\|_{2}^{2}$
21:         $s^{(g)}\leftarrow\log\left(\mathcal{H}(\sigma(f(x)))\cdot\Delta\ell^{(g)}\right)-\log\Delta z^{(g)}$
22:     end for
23:     return $\frac{1}{G}\sum_{g=1}^{G}s^{(g)}$
24:end function
25:
26:function RLT($x$)
27:     $r\leftarrow$ RT($x$),  $l\leftarrow$ LT($x$)
28:     $r_{norm}\leftarrow\Phi^{-1}(\hat{\mathcal{F}}_{RT}(r))$
29:     $l_{norm}\leftarrow\Phi^{-1}(\hat{\mathcal{F}}_{LT}(l))$
30:     return $r_{norm}^{2}+l_{norm}^{2}$
31:end function
In this section, we introduce our layer-wise adversarial example detection framework (Figure 2), comprising three measures: RT, LT, and their fusion RLT. RT captures inconsistencies in intermediate-layer embeddings, while LT measures logit instability under augmented inputs. At test time, given measure $m\in\{\mathrm{RT},\mathrm{LT},\mathrm{RLT}\}\quad\text{and}\quad\tau_{m},$ we classify $\hat{a}(x)=\mathbf{1}^{\top}\bigl{(}f(x)>\tau_{f}\bigr{)}\in\{0,1\}$ as AE detection results. The algorithm for calculating each measure is shown in Algorithm 1. Theoretical assumptions and proofs of our method are shown in Appendix A and Appendix B, respectively.
3.1 Definition and Notations
Given a target DNN $f(\cdot)=f_{logit}\circ f_{L}\circ f_{L-1}\circ\cdots\circ f_{1}(\cdot)$ with $L$ intermediate layers $f_{i}$ and one logit layer $f_{logit}$ at the end, we compute intermediate representations $z_{i}(x_{j})=f_{i}\circ\cdots\circ f_{1}(x_{j})\in\mathbb{R}^{D_{i}}$ for all $i\in\{1,2,\cdots,L\}$, where $D_{i}$ denotes the dimensionality of the embedding space and $x_{j}$ denotes an input sample. We assume the logit layer takes the last intermediate embedding $z_{L}$ to produce the final logit output $\ell=f_{logit}(z_{L})\in\mathbb{R}^{C}$, where $C$ is the number of output classes. We denote the softmax function as $\sigma(x)=\frac{\exp{x}}{\mathbf{1}^{\top}\exp{x}}$, the Shannon entropy as $\mathcal{H}(p)=-\sum p_{i}\log p_{i}$, and the one-hot vector of the ground-truth label $y$ as $\mathbf{o}_{y}\in\{0,1\}^{C}$. Throughout, all auxiliary modules are trained using only benign samples $\mathcal{D}_{norm}=\{x_{1},x_{2},\cdots,x_{N}\}$ without using labels.
3.2 Core Assumption
We begin by formalizing the central structural assumption underlying our detection framework:
Assumption 1 (A Few Large Shifts).
Let $x^{\mathrm{adv}}=x+\delta$ be an adversarial input, and let $z_{k}(x)$ denote the output of $k$-th layer $f_{k}$. We assume there exists a small subset of layers $\mathcal{T}\subseteq\{f_{1},\dots,f_{L},f_{logit}\}$ such that for any $f_{k}\in\mathcal{T}$,
$\frac{1}{D_{k+1}}\|z_{k+1}(x^{\mathrm{adv}})-z_{k+1}(x)\|_{2}^{2}\gg\frac{1}{D_{k}}\|z_{k}(x^{\mathrm{adv}})-z_{k}(x)\|_{2}^{2}.$
For all other layers $f_{j}\notin\mathcal{T}$, the shift between representations remains small:
$\frac{1}{D_{j+1}}\|z_{j+1}(x^{\mathrm{adv}})-z_{j+1}(x)\|_{2}^{2}\approx\frac{1}{D_{j}}\|z_{j}(x^{\mathrm{adv}})-z_{j}(x)\|_{2}^{2}.$
This assumption formalizes the intuition that adversarial perturbations cause disproportionately large changes in representation between specific adjacent layers, while most other transitions remain relatively stable. These localized disruptions manifest as sharp deviations in feature space at only a few critical points in the network. In the following sections, we introduce our proposed testing measures designed to capture and quantify these irregular transitions.
3.3 Recovery Testing Measure
For each hidden layer $k\in\{\,k_{RT},\,k_{RT}+1,\,\ldots,\,L-1\}$, we train an approximate inverse function: $R^{(L\rightarrow k)}:\mathbb{R}^{D_{L}}\longrightarrow\mathbb{R}^{D_{k}}$ to reconstruct $z_{k}$ from $z_{L}$ with hyperparameter $k_{RT}$. Each $R^{(L\rightarrow k)}$ is implemented as a lightweight MLP with 3–4 layers and trained by a mean squared error loss:
$\mathcal{L}_{RT}=\frac{1}{N}\sum_{n=1}^{N}\sum_{k=k_{RT}}^{L-1}\|\,z_{k}(x_{n})-R^{(L\rightarrow k)}\bigl{(}z_{L}(x_{n})\bigr{)}\|_{2}^{2}.$ (1)
At test time, we define the normalized squared error for each layer $k$ as
$e_{k}(x)$ $=\|\,z_{k}(x)-R^{(L\rightarrow k)}(z_{L}(x))\|_{2}^{2},(k=1,\ldots,L-1.)$ (2)
Collecting the vector $e(x)=(e_{1}(x),\ldots,e_{L-1}(x))$, we normalize it with softmax to obtain a distribution $\sigma(e(x))$ across layers. The non-uniformity of this distribution is captured using inverse entropy $\log(L-1)-\mathcal{H}(\sigma(e(x)))$. The final score is computed by taking the average of the raw reconstruction errors and weighting it by this information-based term:
$RT(x)$ $=\bigl{(}\log(L-1)-\mathcal{H}(\sigma(e(x)))\bigr{)}\log\Bigl{(}\frac{1}{L-k_{RT}}\sum_{k=k_{RT}}^{L-1}e_{k}(x)\Bigr{)}.$ (3)
We apply a logarithm to the average of raw errors for numerical stability. Under our assumption, adversarial samples that target specific layers $\mathcal{T}\subseteq\{f_{1},\ldots,f_{L-1}\}$ will yield higher $RT(x)$ values due to large reconstruction errors (see Theorem 1) and a sharply peaked error distribution. Intuitively, the first term of Equation 3 measures how deterministic the error distribution is across layers, while the second term captures the magnitude of a few large shifts.
3.4 Logit-layer Testing Measure
Although RT effectively detects perturbations in intermediate layers $f_{i}$, it cannot be applied to the final logit layer $f_{\text{logit}}$, since there is no subsequent representation from which to reconstruct the logits. To address this—and motivated by the intuition that adversarial examples aim to induce misclassification—we introduce Logit-layer Testing (LT): we quantify uncertainty at $f_{\text{logit}}$, relative to changes in intermediate features, using data-driven, low-cost input augmentations.
Let $\{W^{(g)}\}_{g=1}^{G}$ be a small set (e.g., $1\leq G\leq 6$) of image transformation matrices. Each $W^{(g)}$ is fine-tuned on benign data by minimizing the difference in prediction results, i.e., $\sigma(f(x))\approx\sigma(f(W^{(g)}x))$, while inducing larger perturbations in the intermediate layers. This encourages finding $W^{(g)}$ such that $f_{logit}$ remains robust in its prediction confidence, even when the intermediate layers $f_{i}$ are less robust for benign inputs.
Given a test input, for each augmentation $g$, we measure two types of inconsistencies between the original and augmented inputs: feature-space drift $\Delta z^{(g)}(x|W^{(g)})$ and change in logit decidedness $\Delta\ell^{(g)}(x|W^{(g)})$. The feature-space drift is measured by averaging the $L_{2}$ distances across all intermediate outputs: $\Delta z^{(g)}(x|W^{(g)})=\frac{1}{L-k_{LT}+1}\sum_{i=k_{LT}}^{L}\|z_{i}(x)-z_{i}(W^{(g)}x)\|_{2}^{2}$ with hyperparameter $k_{LT}$. The logit decidedness is measured by computing the $L_{2}$ distance between the predicted one-hot vector and the softmax output: $\Delta\ell^{(g)}(x|W^{(g)})=\|\mathbf{o}_{\hat{y}}-\sigma(f(W^{(g)}x))\|_{2}^{2}$, where $\hat{y}=\arg\max_{c}\sigma_{c}(x)$. We then combine the two quantities into $\frac{\Delta\ell^{(g)}(x|W^{(g)})}{\Delta z^{(g)}(x|W^{(g)})}$, weight it by the entropy of the logit score vector $\mathcal{H}(\sigma(\ell(x)))$ to emphasize inputs closer to the decision boundary [12], and finally average over all augmentations as follows:
$LT(x)=\frac{1}{G}\sum_{g=1}^{G}{\log{\mathcal{H}(\sigma(\ell(x)))\Delta\ell^{(g)}(x|W^{(g)})}-\log{\Delta z^{(g)}(x|W^{(g)})}}.$ (4)
As before, the logarithm ensures numerical stability. Intuitively, under a successful adversarial attack the logit shift $\Delta\ell^{(g)}$ becomes disproportionately large compared to the accumulated feature drift $\Delta z^{(g)}$, which is precisely what LT is designed to detect (Theorem 3).
To fine-tune the transformation matrices $\{W^{(g)}\}$, we directly use $LT(x)$ as the training loss:
$\mathcal{L}_{LT}=\frac{1}{N}\sum_{n=1}^{N}LT(x_{n}).$ (5)
3.5 Recovery and Logit Testing Combined Measure
To capture inconsistencies in both intermediate and logit layers, we introduce a combined score called Recovery and Logit Testing. This score integrates RT and LT while correcting for differences in their statistical distributions using quantile normalization. Specifically, we transform each score based on its empirical cumulative distribution estimated from the benign training set, and map it to the standard normal distribution.
Let ^$\hat{\mathcal{F}}_{RT}$ and ^$\hat{\mathcal{F}}_{LT}$ be the empirical cumulative distribution functions (CDFs) of RT and LT scores computed over the benign data. Let $\Phi^{-1}$ denote the quantile function (inverse CDF) of the standard normal distribution. At test time, each score is transformed as follows:
$RT_{norm}(x)=\Phi^{-1}\left(\hat{\mathcal{F}}_{RT}(RT(x))\right),\quad LT_{norm}(x)=\Phi^{-1}\left(\hat{\mathcal{F}}_{LT}(LT(x))\right).$
The final RLT score is computed by summing the squared normalized values:
$RLT(x)=(RT_{norm}(x))^{2}+(LT_{norm}(x))^{2}.$ (6)
This quantile-based transformation aligns both RT and LT scores to a common standard normal distribution, ensuring that the final combined score reflects significant deviations under either test (Theorem 4), independent of their original scales. This enhances robustness to varied attack types and scoring dynamics.
4 Experiments
To evaluate the effectiveness of our proposed methods, we conduct five main experiments: standard adversarial detection on CIFAR-10 and ImageNet (subsection 4.2); robustness evaluation under adaptive attacks (Orthogonal-PGD) (subsection 4.3); empirical validation of our core assumption on internal feature shifts (subsection 4.4); analysis of computational overhead for efficiency (subsection 4.2); and an ablation study to assess the contribution of each component (Appendix C).
4.1 Experimental Setup
Datasets. We evaluate our method on both CIFAR-10 and ImageNet datasets. CIFAR-10 consists of $60{,}000$ $32\times 32$ images across 10 classes, with standard training (50k) and test (10k) splits. For ImageNet, we use the official training and validation sets, resizing all images to $256\times 256$ and applying standard normalization. CIFAR-10 is used to validate the generality and scalability of our detection approach in a lower-resolution setting, while ImageNet serves as a large-scale benchmark.
Backbone Models. For CIFAR-10, we use a ResNet-110 pretrained classifier. For ImageNet, we employ a pretrained DenseNet-121. In both cases, we extract intermediate features from multiple layers without fine-tuning the backbone, ensuring a consistent foundation for detection.
Threat Models. We evaluate our detection framework under two standard adversarial settings: Limited Knowledge and Perfect Knowledge, following the protocol of previous works [3, 39]. In the Limited Knowledge setting, the adversary has full access to the target classifier but is unaware of the detection mechanism, which remains confidential. In contrast, the Perfect Knowledge (adaptive attack) setting assumes that the adversary has full knowledge of both the classifier and the detection strategy, enabling it to craft attacks specifically to evade detection.
Attack Methods. We evaluate robustness and detection performance under a diverse suite of white-box adversarial attacks, including Fast Gradient Sign Method (FGSM), Projected Gradient Descent (PGD), and Carlini–Wagner (CW) attacks. We also apply AutoAttack (standard version) and Square Attack [2], both of which are parameter-free (black-box) and ensemble-based. To assess robustness under this stronger threat model, we incorporate the state-of-the-art adaptive attack, Orthogonal-PGD [4], which optimizes adversarial perturbations not only to fool the classifier but also to minimize detection signals. This provides a rigorous evaluation of the proposed methods under any conditions.
Implementation Details. All models are implemented in PyTorch and trained on NVIDIA RTX 6000 Ada. We use a batch size of 32, AdamW optimizer with learning rate $1\times 10^{-4}$ and weight decay of $0.01$, training the recover modules and augmentation matrices for 50 epochs.
4.2 Detection Performance under Standard Attacks
Table 2: The AUC of Different Adversarial Detection Approaches on CIFAR-10. The results are the mean and standard deviation of 5 runs. Our methods are included for comparison. Classifier: ResNet110, FGSM: $\epsilon=0.05$, PGD: $\epsilon=0.02$. Note that our methods and BEYOND need no AE for training, leading to the same value on both seen and unseen settings. The bolded values are the best performance, and the underlined italicized values are the second-best performance.
Method Unseen (Attacks in training are excluded from tests) Seen (Attacks in training are included in tests)
FGSM PGD AutoAttack Square FGSM PGD CW AutoAttack Square
DkNN [27] 61.55 ±0.023 51.22 ±0.026 52.12 ±0.023 59.46 ±0.022 61.55 ±0.023 51.22 ±0.026 61.52 ±0.028 52.12 ±0.023 59.46 ±0.022
kNN [9] 61.83 ±0.018 54.52 ±0.022 52.67 ±0.022 73.39 ±0.020 61.83 ±0.018 54.52 ±0.022 62.23 ±0.019 52.67 ±0.022 73.39 ±0.020
LID [23] 71.08 ±0.024 61.33 ±0.025 55.56 ±0.021 66.18 ±0.025 73.61 ±0.020 67.98 ±0.020 55.68 ±0.021 56.33 ±0.024 85.94 ±0.018
Hu [14] 84.51 ±0.025 58.59 ±0.028 53.55 ±0.029 95.82 ±0.020 84.51 ±0.025 58.59 ±0.028 91.02 ±0.022 53.55 ±0.029 95.82 ±0.020
Mao [24] 95.33 ±0.012 82.61 ±0.016 81.95 ±0.020 85.76 ±0.019 95.33 ±0.012 82.61 ±0.016 83.10 ±0.018 81.95 ±0.020 85.76 ±0.019
LNG [1] 98.51 63.14 58.47 94.71 99.88 91.39 89.74 84.03 98.82
BEYOND [39] 98.89 ±0.013 99.28 ±0.020 99.16 ±0.021 99.27 ±0.016 98.89 ±0.013 99.28 ±0.020 99.20 ±0.008 99.16 ±0.021 99.27 ±0.016
Our Approaches
RT 99.93 ±0.005 96.89 ±0.071 $9$0.000 85.38 ±0.344 99.93 ±0.005 96.89 ±0.071 $9$0.002 $9$0.000 85.38 ±0.344
LT 97.50 ±0.038 98.61 ±0.042 99.60 ±0.018 97.47 ±0.036 97.50 ±0.038 98.61 ±0.042 97.08 ±0.027 99.60 ±0.018 97.47 ±0.036
RLT 99.85 ±0.005 99.37 ±0.011 $9$0.000 95.95 ±0.102 99.85 ±0.005 99.37 ±0.011 99.91 ±0.004 $9$0.000 95.95 ±0.102
Table 3: The AUC of Different Adversarial Detection Approaches on ImageNet. To align with baselines, classifier: DenseNet121, FGSM: $\epsilon=0.05$, PGD: $\epsilon=0.02$. Due to memory and resource constraints, baseline methods are not evaluated against AutoAttack on ImageNet.
Method Unseen Seen
FGSM PGD FGSM PGD CW
DkNN [27] 89.16 ±0.038 78.00 ±0.041 89.16 ±0.038 78.00 ±0.041 68.91 ±0.044
kNN [9] 51.63 ±0.04 51.14 ±0.039 51.63 ±0.04 51.14 ±0.039 50.73 ±0.04
LID [23] 90.32 ±0.046 52.56 ±0.038 99.24 ±0.043 98.09 ±0.042 58.83 ±0.041
Hu [14] 72.56 ±0.037 86.00 ±0.042 72.56 ±0.037 86.00 ±0.042 80.79 ±0.044
LNG [1] 96.85 89.61 99.53 98.42 86.05
BEYOND [39] 97.59 ±0.04 96.26 ±0.045 97.59 ±0.04 96.26 ±0.045 95.46 ±0.047
Our Approaches
RT 94.31 ±0.457 $9$0.000 94.31 ±0.457 $9$0.000 92.18 ±0.135
LT 96.18 ±0.028 97.89 ±0.021 96.18 ±0.028 97.89 ±0.021 94.06 ±0.215
RLT 97.60 ±0.048 $9$0.000 97.60 ±0.048 $9$0.000 91.19 ±0.022
We evaluate the effectiveness of our proposed detection scores, namely RT, LT, and RLT, on CIFAR-10 and ImageNet under standard adversarial threat models, including FGSM, PGD, CW, AutoAttack, and Square Attack. We benchmark against established baselines such as LID [23], DkNN [27], LNG [1], and the recent SSL-based BEYOND [39]. The Area Under the Receiver Operating Characteristic Curve (RoC-AUC) results, which measure detection performance on various thresholds, are reported in Table 2 for CIFAR-10 and Table 3 for ImageNet.
On CIFAR-10, RT achieves exceptionally strong performance across attacks that disrupt internal representations (e.g., FGSM, CW, AutoAttack), aligning with its role in capturing large, localized deviations in intermediate layers. LT performs particularly well against attacks like PGD and Square, which introduce minimal internal distortion but induce instability at the output layer. Notably, the fused RLT score achieves either the best or second-best AUC across most of the attacks, validating the complementary nature of RT and LT. Despite relying solely on the internal signals of the target classifier and requiring no adversarial examples or external models, RLT matches or outperforms state-of-the-art methods, including BEYOND, which uses large pre-trained SSL representations.
On ImageNet, we observe similar trends. RT continues to perform well under strong perturbations such as PGD, while LT maintains robust accuracy under less structured noise like FGSM. The combined RLT score again leads to superior or competitive detection performance. Unlike several baselines that degrade substantially in the large-scale setting due to increased model capacity or overfitting to specific attacks, our method remains stable without requiring additional training resources or architecture-specific tuning.
4.3 Detection Performance under Adaptive Attacks
To assess the robustness of our proposed method against strong adaptive adversaries, we evaluate the RLT score under the Orthogonal-PGD adaptive attack, following the protocol in BEYOND [39], where the attacker is assumed to have full access to both the target classifier and the detection mechanism, and crafts perturbations specifically to evade detection by inducing misclassification.
Table 4 reports the robust accuracy (RA) of various defense methods at two $L_{\infty}$ threat levels ($0.01$ and $8/255$), each evaluated at false positive rates (FPR) of 5% and 50%. While BEYOND achieves strong performance with the aid of additional components such as self-supervised models, our method achieves competitive performance using only the internal consistency of the target model, without requiring adversarial examples or external models during training. These results highlight the effectiveness of our RLT-based detection framework even in adversarially optimized threat settings.
Table 4: RA(%) under Orthogonal-PGD Adaptive Attack using CIFAR-10 and ResNet110.
Defense $\bm{L_{\infty}=0.01}$ $\bm{L_{\infty}=8/255}$
RA@FPR5% RA@FPR50% RA@FPR5% RA@FPR50%
Ours (RLT) 75.40 99.58 33.70 80.77
BEYOND [39] 88.38 98.81 13.80 48.20
Trapdoor [29] 0.00 7.00 0.00 8.00
DLA [31] 62.60 83.70 0.00 28.20
SID [32] 6.90 23.40 0.00 1.60
SPAM [19] 1.20 46.00 0.00 38.00
4.4 Empirical Evaluation of the Proposed Assumption
Figure 3: Empirical validation of the A Few Large Shifts Assumption using CIFAR-10 and ResNet-110. We plot layer-wise error distributions $\sigma(e_{k})$ under different attack methods.
To validate our A Few Large Shifts Assumption, we measured layer-wise reconstruction error distributions $\sigma(e_{k})$ on a ResNet-110 model trained on CIFAR-10 under various attacks: FGSM, PGD, CW, AutoAttack, and Square Attack. As shown in Figure 3, benign inputs exhibit relatively flat error profiles across layers, while certain attacks as shown in Figure 3.(a), produce sharp peaks at specific layers—indicating large indicated as red arrows, localized shifts in internal representations. These peaked patterns align with high RT detection performance and support the assumption that adversarial perturbations affect only a subset of layers $\mathcal{T}\subseteq\{f_{1},f_{2},\cdots,f_{L}\}$. In contrast, attacks like PGD and Square, which show stronger LT detection, yield smoother error curves similar to benign inputs but with greater variance as depicted in Figure 3.(b). This suggests that these attacks primarily affect the logit layer, with minimal disruption to intermediate features $\mathcal{T}=\{f_{logit}\}$. These findings empirically confirm that adversarial manipulations often concentrate at a few layers and further justify the complementary roles of RT and LT in our detection framework.
4.5 Implementation Costs
We assess the computational efficiency of our RLT method against existing baselines. As shown in Table 5, RLT introduces minimal floating-point operations (FLOPs), parameters, and model size overhead, offering a lightweight, plug-in solution without retraining or external models. It is significantly more efficient than SSL-based (Mao et al.[24]) and graph-based (LNG[1]) detectors, making it well-suited for real-time and resource-constrained deployments.
Table 5: Comparison of implementation costs in terms of FLOPs, parameters, and model size overhead when the target model is ResNet110 with the CIFAR-10 dataset.
Method FLOPs (G) Params (M) Model Size Overhead ($\bm{\times}$)
Mao [24] 5.25 38.12 22.02
LNG [1] 0.286 8.33 4.81
BEYOND [39] 0.715 20.62 11.91
Ours (RLT) 0.491 2.59 1.49
5 Limitations and Future Work
Our proposed framework uses RT and LT to detect adversarial examples by leveraging internal layer inconsistencies. Despite its simplicity, efficiency, and strong performance across diverse settings, it has several limitations that we now discuss.
Our framework hinges on the A Few Large Shifts principle—that adversarial perturbations manifest predominantly in either a small subset of internal layers $\{f_{1},\dots,f_{L}\}$ or (separately) at the logit stage. While this binary decomposition keeps RT and LT analytically clean and easy to interpret, in practice highly adaptive or high-capacity attacks may simultaneously disrupt multiple layers and the logit outputs, creating overlapping distortions. In such mixed-mode scenarios, our current score fusion may underestimate these interactions, potentially reducing detection accuracy. Evaluating the detector under attacks designed to violate this assumption—and developing extensions to handle truly mixed shifts—constitutes important future work.
Moreover, the logit space itself poses unique challenges for detection. Unlike intermediate feature maps, which admit spatial or structural comparisons, logits lack an interpretable geometric structure and offer no downstream transform against which to gauge stability. As a result, it is difficult to distinguish between benign variability—due to natural input perturbations or augmentations—and adversarial influence. Although our LT score partially mitigates this by using entropy-weighted changes under augmentations, it remains only a coarse proxy for true perturbation sensitivity.
Finally, following prior work [39], we have evaluated our method only on CNNs (e.g., ResNet-110) and vision benchmarks. Extending our approach to modern architectures such as Transformers [33], as well as stress-testing it on assumption-violating and mixed-mode attacks, are important directions for future work.
6 Conclusion
In this work, we introduced a novel framework for adversarial example detection that leverages layer-wise inconsistencies within deep neural networks. Motivated by our proposed A Few Large Shifts Assumption–which posits that adversarial perturbations primarily cause substantial deviations in a small subset of network layers–we developed two complementary detection strategies: RT, capturing discrepancies in intermediate-layer reconstructions, and LT, quantifying output-layer sensitivity under minimal augmentations. Extensive evaluations on CIFAR-10, CIFAR-100, and ImageNet demonstrate that our combined approach, RLT achieves state-of-the-art detection performance, significantly outperforming existing baselines while remaining highly robust to adaptive adversarial attacks. Our method operates efficiently without relying on external models, adversarial training, or extensive augmentation, highlighting its practical suitability for real-world deployment.
Acknowledgements
This work was supported in part by the DARPA Young Faculty Award, the National Science Foundation (NSF) under Grants #2127780, #2319198, #2321840, #2312517, and #2235472, the Semiconductor Research Corporation (SRC), the Office of Naval Research through the Young Investigator Program Award, and Grants #N00014-21-1-2225 and #N00014-22-1-2067, Army Research Office Grant #W911NF2410360, and the U.S. Army Combat Capabilities Development Command (DEVCOM) Army Research Laboratory under Support Agreement No. USMA 21050. Additionally, support was provided by the Air Force Office of Scientific Research under Award #FA9550-22-1-0253, along with generous gifts from Xilinx and Cisco. The views expressed in this paper are those of the authors and do not reflect the official policy or position of the U.S. Military Academy, the U.S. Army, the U.S. Department of Defense, or the U.S. Government.
Appendix A Theoretical Analysis
We provide formal justification for our proposed detection framework, which includes RT, LT, and the fused score RLT. For each score, we analyze its behavior under benign and adversarial inputs, showing that the design of the metric leads to statistically separable score distributions. This justifies the detection reliability of our method under common threat models.
A.1 Justification for Layer-wise Reconstruction Error in RT
This section analyzes why using the reconstruction error of intermediate features—as done in RT—produces a separable distribution between benign and adversarial inputs. Specifically, we show that benign inputs produce consistently low recovery residuals, while adversarial perturbations, though small in input space, induce disproportionately large deviations in the representation space at select layers, leading to higher RT scores.
Assumption 2 (Approximate Invertibility).
For each intermediate layer $i\in\{k_{RT},\dots,L-1\}$, there exists a well-trained inverse function $R^{(L\rightarrow i)}$ such that, for benign inputs $x$,
$\|z_{i}(x)-R^{(L\rightarrow i)}(z_{L}(x))\|\leq\varepsilon,$
for some small constant $\varepsilon>0$.
Assumption 3 (Flatness of the Inverse Function).
The inverse function $R^{(L\rightarrow i)}$ is Lipschitz-smooth with constant $\alpha\ll 1$. That is, for all small perturbations $\delta$,
$\|R^{(L\rightarrow i)}(z_{L}(x+\delta))-R^{(L\rightarrow i)}(z_{L}(x))\|\leq\alpha\cdot\|z_{L}(x+\delta)-z_{L}(x)\|.$
Assumption 4 (Sub-Gaussian Layer Perturbations).
Let $\delta$ be an adversarial perturbation. Then the induced change in intermediate features is sub-Gaussian:
$\|\nabla z_{i}(x)\cdot\delta\|\sim\text{SubG}(\mu,\sigma^{2}),$
i.e., for all $t>0$,
$\Pr\left(\|\nabla z_{i}(x)\cdot\delta\|\leq\mu-t\right)\leq\exp\left(-\frac{t^{2}}{2\sigma^{2}}\right).$
Theorem 1 (RT Detects Adversarial Residuals).
Under Assumption 1, we assume that $\mathcal{T}\subseteq\{f_{1},\dots,f_{L-1}\}$ contains one or more intermediate layers where adversarial perturbations induce disproportionately large shifts. These shifts lead to elevated reconstruction residuals, which RT is designed to detect. Let $x^{\mathrm{adv}}=x+\delta$ be an adversarial example, and suppose Assumption 2, 3, and 4 hold. Then, with probability at least $1-\eta$ for small $\eta$, the following inequality holds:
$\|z_{i}(x^{\mathrm{adv}})-R^{(L\rightarrow i)}(z_{L}(x^{\mathrm{adv}}))\|^{2}>\|z_{i}(x)-R^{(L\rightarrow i)}(z_{L}(x))\|^{2}.$
(Proof: Appendix B.1)
A.2 Justification for Ratio-based Logit-layer Deviation in LT
In this section, we justify the use of the LT score, which captures how the logit-layer output changes under augmentation relative to changes in intermediate-layer features. We show that benign inputs maintain logit consistency under mild perturbations, while adversarial examples—particularly those crafted to flip decisions—exhibit exaggerated logit volatility, causing the LT score to grow disproportionately.
Assumption 5 (Benign Augmentation Stability).
For any benign input $x$, and a mild transformation $W^{(g)}$, the resulting features satisfy
$\|z^{(g)}-z\|\leq\eta,\quad\|\ell^{(g)}-\ell\|\leq\alpha(\eta),$
for small $\eta>0$ and monotonically increasing $\alpha(\eta)\ll 1$.
Assumption 6 (Recovery Test Evasion).
Let $x^{\mathrm{adv}}=x+\delta$ be an adversarial input that satisfies Assumption 1 with $\mathcal{T}=\{f_{logit}\}$, i.e., the perturbation induces a large shift only at the logit layer, while intermediate representations remain largely consistent with those of benign inputs.
Then the residuals measured by RT remain low:
$\|z_{i}(x^{\mathrm{adv}})-R^{(L\rightarrow i)}(z_{L}(x^{\mathrm{adv}}))\|^{2}\lesssim\|z_{i}(x)-R^{(L\rightarrow i)}(z_{L}(x))\|^{2},$
yet flips the final prediction:
$\arg\max\ell(x^{\mathrm{adv}})\neq\arg\max\ell(x).$
Moreover, augmentations preserve feature drift:
$\|z^{(g),\text{adv}}-z^{\text{adv}}\|\approx\|z^{(g)}-z\|.$
Theorem 2 (Logit Instability under Augmentation).
Under Assumption 6, the adversarial logit output is unstable under small augmentation:
$\|\textbf{o}(\ell(x^{\mathrm{adv}}))-\textbf{o}(\ell(x^{(g),\text{adv}}))\|>0.$
(Proof: Appendix B.2)
Theorem 3 (Amplified Logit Sensitivity).
Under Assumption 5 and Assumption 6, the logit sensitivity of adversarial inputs satisfies:
$\frac{\|\ell^{(g),\text{adv}}-\ell^{\text{adv}}\|}{\|z^{(g),\text{adv}}-z^{\text{adv}}\|}>\frac{\|\ell^{(g)}-\ell\|}{\|z^{(g)}-z\|}.$
(Proof: Appendix B.3)
A.3 Quantile-normalized RT + LT Provides Jointly Separable Score in RLT
We now provide a theoretical justification for the fused detection score $RLT(x)=RT_{\text{norm}}^{2}+LT_{\text{norm}}^{2}$. Since RT and LT each capture different types of adversarial signatures (internal layers misalignment vs. logit instability), combining them creates a more robust metric. By applying quantile normalization, we map both scores into a common distributional space, ensuring fair fusion. We then show that the fused score statistically separates adversarial inputs even when only one metric is significantly perturbed.
Assumption 7 (Quantile-normalized RT and LT).
Let ^$\hat{\mathcal{F}}_{RT},\hat{\mathcal{F}}_{LT}$ be empirical CDFs computed on benign RT and LT scores. Define:
$RT_{\text{norm}}(x)=\Phi^{-1}(\hat{\mathcal{F}}_{RT}(RT(x))),\quad LT_{\text{norm}}(x)=\Phi^{-1}(\hat{\mathcal{F}}_{LT}(LT(x))),$
where $\Phi^{-1}$ is the standard normal quantile function.
Assumption 8 (Adversarial Score Margin).
There exists $\gamma>0$ such that for adversarial $x$, at least one normalized score satisfies:
$|RT_{\text{norm}}(x)|>\gamma\quad\text{or}\quad|LT_{\text{norm}}(x)|>\gamma.$
Theorem 4 (RLT Separates Adversaries).
Define the fused score:
$RLT(x)=RT_{\text{norm}}(x)^{2}+LT_{\text{norm}}(x)^{2}.$
Then, under Assumption 7 and Assumption 8,
$\mathbb{E}[RLT(x)]=2\quad\text{for benign }x,\quad\text{and}\quad RLT(x^{\text{adv}})>\gamma^{2}.$
(Proof : Appendix B.4)
A.4 Robustness to Adaptive Attacks
In this section, we provide a theoretical perspective on the robustness of our detection framework under adaptive adversaries. An adaptive attack refers to a threat model where the adversary has full knowledge of both the classifier and the detection mechanism, and explicitly optimizes its objective to evade detection. We show that the design of our RT and LT metrics inherently introduces conflicting optimization gradients, which hinder the adversary’s ability to jointly suppress both detection scores.
A.4.1 Adaptive Attack Objective
To simulate a strong white-box adversary, we formulate an adaptive objective that simultaneously minimizes the classification loss and both detection scores (RT and LT). Specifically, the adversary solves the following constrained optimization problem:
$\min_{\|\delta\|\leq\epsilon}\mathcal{L}_{\mathrm{cls}}(x+\delta,y_{t})+\beta_{1}\cdot RT(x+\delta)+\beta_{2}\cdot LT(x+\delta),$
where $\epsilon$ is perturbation strength, $\mathcal{L}_{\mathrm{cls}}$ is the classification loss (e.g., cross-entropy), $y_{t}$ is a target label in the case of targeted attacks, and $\beta_{1},\beta_{2}$ are non-negative coefficients controlling the importance of detection suppression. While the classification term drives the adversarial objective, the additional RT and LT terms introduce competing constraints on internal and output-layer consistency. The optimization becomes increasingly difficult when the gradients corresponding to RT and LT conflict in direction.
A.4.2 Conflicting Gradient Effects
The RT score is defined as:
$RT(x)\propto\|z_{i}(x)-R^{(L\rightarrow i)}(z_{L}(x))\|^{2},$
which penalizes deviations from the inverse-mapped intermediate representations. Minimizing this score encourages the adversary to maintain stable internal features consistent with benign patterns. In contrast, the LT score is defined as:
$LT(x)\propto\frac{\mathcal{H}(\sigma(\ell(x)))\cdot\Delta\ell^{(g)}(x)}{\Delta z^{(g)}(x)},$
where $\Delta\ell^{(g)}(x)$ measures logit deviation across augmentations, and $\Delta z^{(g)}(x)$ captures the corresponding feature drift. Minimizing LT encourages logit stability while allowing some augmentation-induced feature variability. These two goals inherently compete. Enforcing small residuals in intermediate layers (RT) limits the allowable variation in augmented features, which can inflate LT. Conversely, promoting augmentation-invariant logits (LT) often introduces instability in internal features, increasing RT. As a result, gradients of the two terms may partially oppose each other, degrading the effectiveness of adversarial optimization. This conflict acts as a built-in regularizer against adaptive attacks. In practice, we observe that attempting to suppress one score often exacerbates the other, making it difficult to minimize the fused detection score $RLT(x)=RT_{\text{norm}}^{2}(x)+LT_{\text{norm}}^{2}(x)$.
Appendix B Proofs
B.1 Theorem 1
Proof.
For the benign input $x$, Assumption 2 guarantees:
$\|z_{i}(x)-R^{(L\rightarrow i)}(z_{L}(x))\|\leq\varepsilon.$
Now consider the adversarial input $x^{\mathrm{adv}}=x+\delta$. Using first-order Taylor expansions:
$z_{i}(x^{\mathrm{adv}})=z_{i}(x)+\nabla z_{i}(x)\cdot\delta+o(\|\delta\|),$
$z_{L}(x^{\mathrm{adv}})=z_{L}(x)+\nabla z_{L}(x)\cdot\delta+o(\|\delta\|).$
Next, apply the inverse recovery map:
$R^{(L\rightarrow i)}(z_{L}(x^{\mathrm{adv}}))=R^{(L\rightarrow i)}(z_{L}(x))+\nabla R^{(L\rightarrow i)}(z_{L}(x))\cdot(z_{L}(x^{\mathrm{adv}})-z_{L}(x))+o(\|\delta\|).$
Define:
$\Delta_{z}:=\nabla z_{i}(x)\cdot\delta,\quad\Delta_{r}:=\nabla R^{(L\rightarrow i)}(z_{L}(x))\cdot\nabla z_{L}(x)\cdot\delta.$
Then:
$z_{i}(x^{\mathrm{adv}})-R^{(L\rightarrow i)}(z_{L}(x^{\mathrm{adv}}))\approx(z_{i}(x)-R^{(L\rightarrow i)}(z_{L}(x)))+(\Delta_{z}-\Delta_{r}).$
By reverse triangle inequality:
$\|z_{i}(x^{\mathrm{adv}})-R^{(L\rightarrow i)}(z_{L}(x^{\mathrm{adv}}))\|\geq\|\Delta_{z}\|-\|z_{i}(x)-R^{(L\rightarrow i)}(z_{L}(x))\|-\|\Delta_{r}\|.$
From Assumption 2 and Assumption 3:
$\|z_{i}(x)-R^{(L\rightarrow i)}(z_{L}(x))\|\leq\varepsilon,\quad\|\Delta_{r}\|\leq\alpha\cdot\|\nabla z_{L}(x)\cdot\delta\|.$
Assuming $\mu>2(\varepsilon+\alpha\cdot\|\nabla z_{L}(x)\cdot\delta\|)$, Assumption 4 implies that with probability at least $1-\eta$,
$\|\Delta_{z}\|\geq\mu>2\varepsilon+\|\Delta_{r}\|,$
where $\eta\leq\exp\left(-\frac{(\mu-2(\varepsilon+\|\Delta_{r}\|))^{2}}{2\sigma^{2}}\right).$ Then:
$\|z_{i}(x^{\mathrm{adv}})-R^{(L\rightarrow i)}(z_{L}(x^{\mathrm{adv}}))\|\geq\varepsilon.$
Squaring both sides proves the theorem. ∎
B.2 Theorem 2
Proof.
Given the prediction is flipped but the intermediate features remain close to those of $x$, the adversarial logit lies near a decision boundary. Thus, even a mild augmentation $x^{(g),\text{adv}}$ can shift the logits across the boundary, changing the predicted class. ∎
B.3 Theorem 3
Proof.
Assume the contrary. Then adversarial sensitivity is less than benign. Given $\|z^{(g),\text{adv}}-z^{\text{adv}}\|\approx\|z^{(g)}-z\|$, this implies:
$\|\ell^{(g),\text{adv}}-\ell^{\text{adv}}\|\leq\|\ell^{(g)}-\ell\|\leq\alpha(\eta).$
Hence, $\ell^{(g),\text{adv}}\approx\ell^{\text{adv}}$, contradicting Theorem 2. Thus the adversarial logit must be more sensitive to benign augmentation than the original. ∎
B.4 Theorem 4
Proof.
Since both normalized scores follow $\mathcal{N}(0,1)$, the expected value of their squared sum under benign data is:
$\mathbb{E}[RLT(x)]=\mathbb{E}[RT_{\text{norm}}^{2}]+\mathbb{E}[LT_{\text{norm}}^{2}]=1+1=2.$
For adversarial inputs, Assumption 8 ensures at least one squared score exceeds $\gamma^{2}$. Thus,
$RLT(x^{\text{adv}})>\gamma^{2}.$
Appendix C Ablation Study
All ablation studies were conducted using the CIFAR-10 dataset with ResNet110. We systematically vary the key hyperparameters and architectural choices to assess their impact on detection performance.
C.1 Size of recover module
We evaluate the impact of varying the depth and dimensionality of the recover modules on RT detection performance. As we reduce the size of the recover modules in terms of both depth and dimensionality, we observe slightly improved performance as shown in Table 6. This result aligns well with Assumption 3, which posits that a flatter or smoother recovery module—achieved by reducing complexity—results in improved detection due to closer alignment with the assumption of stable layer-wise reconstruction under benign conditions. However, the overall performance differences are relatively minor, demonstrating robustness to the choice of recover module size.
Table 6: RT test AUC (%) differences when varying depth and dimensionality of recover modules, compared to depth = 5 and dimensionality = 512.
Depth FGSM PGD CW AutoAttack Square Avg.
2 +0.55 -2.03 +0.16 0.00 +10.96 +1.93
3 +0.15 +0.15 +0.08 0.00 +4.25 +0.92
4 -0.01 +0.01 +0.04 0.00 -0.50 -0.09
5 0.00 0.00 0.00 0.00 0.00 0.00
Dimensionality FGSM PGD CW AutoAttack Square Avg.
64 -0.50 -0.10 -0.40 0.00 +8.34 +1.47
128 -0.54 -0.04 -0.20 0.00 +4.13 +0.67
256 -0.18 -0.18 -0.03 0.00 +0.65 +0.13
512 0.00 0.00 0.00 0.00 0.00 0.00
C.2 Number of learnable augmentations $G$
We investigate how varying the number of learnable augmentation matrices ($G$) impacts LT detection performance. As depicted in Table 7, the detection performance exhibits minimal sensitivity to the number of augmentations used. Even with a significantly smaller number of augmentations ($G<4$), LT continues to perform robustly. This indicates our LT method effectively quantifies logit-layer perturbations without relying on extensive augmentation, in contrast to previous methods such as BEYOND [39], which required up to 50 augmentations. The minimal requirement of augmentations highlights our method’s computational efficiency and practical deployability.
Table 7: LT test AUC (%) differences when varying the number of augmentation matrices ($G$), compared to $G=4$.
$\bm{G}$ FGSM PGD CW AutoAttack Square Avg.
1 +0.01 -0.03 -0.02 -0.02 +0.01 -0.01
2 -0.04 +0.08 -0.14 +0.37 +0.03 +0.06
3 0.00 +0.20 -0.05 +0.46 +0.02 +0.12
4 0.00 0.00 0.00 0.00 0.00 0.00
C.3 Choice of $k_{RT}$ and $k_{LT}$
We examine how choosing identical values for hyperparameters $k_{RT}$ and $k_{LT}$ affects combined RLT performance relative to optimal, independent selection of these parameters. Results presented in Table 8 indicate a maximum performance decrease of only 1.33% when setting $k=k_{RT}=k_{LT}$ uniformly, as opposed to independently optimizing each hyperparameter. These findings highlight the practical robustness of our detection framework to hyperparameter selection, emphasizing that fine-grained tuning of $k_{RT}$ and $k_{LT}$ is unnecessary for achieving high detection accuracy, greatly simplifying deployment.
Table 8: RLT test AUC (%) differences for varying values of $k=k_{RT}=k_{LT}$ compared to optimal separate selection.
$\bm{k=k_{RT}=k_{LT}}$ FGSM PGD CW AutoAttack Square Avg.
1 -1.59 -3.45 -0.82 -0.03 -0.77 -1.33
10 -1.37 -3.55 -0.49 -0.04 -0.86 -1.26
15 -1.03 -3.41 -0.35 -0.04 -1.02 -1.17
20 -1.25 -3.16 -0.33 -0.04 -0.63 -1.08
25 -1.37 -2.51 -0.37 -0.03 -1.05 -1.06
30 -1.67 -1.33 -0.39 -0.01 -0.77 -0.83
$k_{RT}\neq k_{LT}$ (Optimal) 0.00 0.00 0.00 0.00 0.00 0.00
C.4 Single objective adaptive attacks
Table 9: Single Objective Orthogonal-PGD Adaptive Attack on $L_{\infty}=8/255$ using CIFAR-10 and ResNet110.
Removed Objective RA@FPR5% RA@FPR50%
None 33.70% 80.77%
RT 17.11% 55.30%
LT 17.20% 56.17%
To further elucidate the inherent robustness of our detection framework, we conduct an ablation analysis by removing individual objectives from the Orthogonal-PGD adaptive attack. Specifically, we investigate scenarios where only one detection objective–either RT or LT–is targeted by the adversary. The results, summarized in Table 9, highlight significant performance drops when either RT or LT objectives are excluded. This empirical outcome clearly substantiates our theoretical analysis presented earlier: adaptive attacks face inherent difficulties due to conflicting gradient directions induced by RT and LT, which makes attack algorithms difficult to find $\mathcal{T}$.
The interpretation of this phenomenon aligns precisely with our conceptual framework. When RT is removed, adversarial perturbations predominantly shift their targeted layers $\mathcal{T}$ towards intermediate layers, significantly affecting internal representations. Conversely, when LT is excluded, adversarial perturbations primarily influence the logit layer, creating instability at the output level. These distinct perturbation strategies confirm our assumption that RT and LT effectively impose opposing constraints on adaptive adversaries, complicating their optimization efforts to simultaneously minimize both scores.
Remarkably, even under this weakened adaptive setting–where adversaries optimize only against one detection objective–our combined detection (RLT) continues to outperform current state-of-the-art baseline models, as reported in Table 4. This robust performance demonstrates that our quantification of layer-by-layer perturbations effectively captures adversarial manipulation, affirming the design efficacy of our proposed RT and LT measures in practical adversarial settings.
C.5 Ablation on perturbation budget sensitivity of each testing measure
Figure 4: Ablation study showing AUC (%) of RT, LT, and RLT under varying perturbation budgets ($\epsilon\in\{0.01,0.02,\cdots,0.1\}$) across multiple standard attack types (FGSM, PGD, AutoAttack, Square). Evaluated on CIFAR-10 using a ResNet-110 classifier and detection models trained with fewer epochs for efficiency.
To analyze the contribution of each detection component across adversarial strengths, we conduct an ablation study varying the perturbation budget $\epsilon$ for common standard (non-adaptive) attacks, including FGSM, PGD, AutoAttack, and Square. We report AUC detection performance for RT, LT, and RLT. All experiments are performed on the CIFAR-10 dataset using a pretrained ResNet-110 classifier. Detection models are trained with a reduced number of epochs to simulate lightweight deployment.
As visualized in Figure 4, detection performance consistently improves as the adversary’s perturbation budget increases. This trend reflects our framework’s strength in leveraging internal inconsistencies that become more pronounced under stronger attacks. Unlike previous methods, such as BEYOND [39], which tend to suffer degraded performance at higher perturbation levels due to their reliance on final feature stability, our method benefits from capturing the greater representational disruption induced across the layers by adversaries at higher budgets.
We also observe that as the budget increases, the targeted layers by adversarial perturbations tend to shift from the logit layer to deeper intermediate layers. This is evidenced by the decline in LT’s AUC and the concurrent improvement in RT’s performance. We interpret this behavior as a reflection of the limited perturbation capacity of each layer, under the assumption that each layer contains only a finite set of vulnerable manifolds. As the perturbation budget grows, the adversary exhausts the capacity of the logit layer and is forced to exploit deeper, intermediate representations. Despite this divergence, RLT–which combines both signals–demonstrates a smooth and consistent increase in performance, validating its robustness and complementary design.
This supports our A Few Large Shifts Assumption, reinforcing that adversarial perturbations typically cause disproportionately large disruptions in a small subset of layers, which become more detectable as the attack budget increases. The results further demonstrate that combining RT and LT provides stable and effective detection across a wide perturbation spectrum.
C.6 Contribution of each term in testing measures
Figure 5: Empirical distributions of individual terms used in our detection metrics, evaluated on CIFAR-10 with a ResNet-110 model under FGSM attack ($\epsilon=0.05$). Scores are squared and normalized using quantile normalization based on clean test samples.
Table 10: Contribution of each term in the LT and RT metrics. AUC scores (in percentage) are reported on CIFAR-10 using ResNet-110 under different attacks. The last column indicates the average drop in performance from the full model.
Removed Term FGSM PGD CW AutoAttack Square Average Drop
Logit-layer Testing (LT)
None (Full LT) 97.50 98.61 97.08 99.60 97.47 98.05
   – Logit Decidedness 99.51 94.66 82.97 96.42 95.91 93.89 -4.16
   – Logit Entropy 97.51 59.37 89.76 57.14 97.20 80.19 -17.86
   – Feature Drift 94.62 65.49 99.29 50.01 97.29 81.34 -16.71
Recovery Testing (RT)
None (Full RT) 99.93 96.89 99.90 99.99 85.38 96.42
   – Inverse Entropy 98.98 94.35 99.09 99.95 84.52 95.38 -1.04
   – Log Error 80.34 97.09 88.95 99.95 65.68 86.40 -10.02
To better understand the role of each component in our proposed detection metrics, we perform an ablation study by individually removing key terms from the Logit-layer Testing and Recovery Testing scores. We evaluate the resulting change in detection performance using AUC scores under five adversarial attack types–FGSM, PGD, CW, AutoAttack, and Square–on CIFAR-10 with a ResNet-110 classifier. Table 10 summarizes the results along with the average performance drop caused by each ablation.
We observe that LT is particularly sensitive to the removal of the logit entropy and feature drift terms. Excluding either leads to a substantial degradation in performance across all attacks, suggesting that these terms are central to LT’s ability to quantify abnormal sensitivity in the output layer relative to internal representation changes. The entropy term, in particular, governs the calibration of the detector based on output uncertainty and appears essential for capturing confidence shifts introduced by adversarial perturbations. The feature drift term complements this by providing a baseline of expected internal variation under benign augmentations. Removing the logit decidedness term results in more modest, yet consistent, performance degradation, indicating its auxiliary role in sharpening the detection boundary.
In the case of RT, the core driver of performance is the reconstruction error between predicted and actual intermediate features. Its removal yields a pronounced drop in detection accuracy, reaffirming that adversarial perturbations often manifest as sharp deviations in the internal feature manifold. In contrast, the inverse entropy term, which weights errors based on the sharpness of their distribution across layers, contributes more marginally to the overall score. Its absence slightly affects performance, implying it primarily serves as a regularizer rather than a principal signal.
These findings underscore that both RT and LT derive their strength from distinct but complementary cues–internal feature consistency and output–level robustness, respectively. While some components act as critical discriminators, others enhance calibration and stability. Together, their integration within the full detection framework leads to robust and generalizable performance across diverse attack types.
Figure 5 visually complements these results by showing the empirical distributions of each term on clean and adversarial samples. We observe that each term, when considered independently, introduces a measurable degree of separation between normal and adversarial inputs. This separation is modest for individual components but becomes substantially more pronounced when the terms are combined within LT and RT. The final integration of both methods in RLT leads to the strongest separation, illustrating the complementary nature of these features.
Appendix D AEs Detection Across Architectures on CIFAR-100
Table 11: AUC scores (%) on CIFAR-100 under three white-box attacks (FGSM, $\ell_{\infty}$-PGD, and AutoAttack) with the same perturbation budget as the previous evaluations.
ResNet-18 MobileNet-V2 x0 5
Variant FGSM PGD AutoAttack Avg. FGSM PGD AutoAttack Avg.
RT 87.83 99.21 99.98 95.67 99.02 99.31 99.99 99.44
LT 97.83 97.53 99.47 98.28 94.73 87.50 91.74 91.33
RLT 95.61 98.92 99.97 98.17 98.73 98.14 99.98 98.95
ShuffleNet-V2 x0 5 RepVGG a0
RT 98.38 97.35 99.97 98.57 91.77 98.94 99.99 96.90
LT 94.82 90.34 96.81 93.99 90.38 96.57 99.26 95.40
RLT 98.16 96.17 99.94 98.09 93.49 98.46 99.98 97.31
To assess the generalizability of our detection framework across diverse neural architectures, we evaluate RT, LT, and RLT on CIFAR-100 using four distinct backbone models: ResNet-18, MobileNet-V2 x0 5, ShuffleNet-V2 x0 5, and RepVGG-a0. As shown in Table 11, our methods consistently achieve high AUC scores across all architectures and attack types (FGSM, PGD, and AutoAttack), with RLT offering the most robust overall performance.
These results demonstrate that our detection strategies—rooted in internal layer inconsistency—are not only effective on traditional architectures like ResNet and DenseNet, but also extend well to lightweight and mobile-friendly models such as MobileNet and ShuffleNet, as well as convolutional variants like RepVGG. Notably, the fused score RLT maintains average AUCs above 97% on all target models, confirming the adaptability and resilience of our approach.
Importantly, the deployment of our detectors requires minimal architectural modification and no adversarial training, making them easily integrable into a wide range of existing models. This plug-and-play nature, combined with high detection efficacy, makes our method highly suitable for practical applications where architectural diversity and efficiency are key considerations.

### v5

@@
 Abstract
-[…]
+Deep neural networks (DNNs) are highly susceptible to adversarial examples—subtle, imperceptible perturbations that can lead to incorrect predictions.
+While detection-based defenses offer a practical alternative to adversarial training, many existing methods depend on external models, complex architectures, or adversarial data, limiting their efficiency and generalizability.
+We introduce a lightweight, plug-in detection framework that leverages internal layer-wise inconsistencies within the target model itself, requiring only benign data for calibration.
+Our approach is grounded in the A Few Large Shifts Assumption, which posits that adversarial perturbations induce large, localized violations of layer-wise Lipschitz continuity in a small subset of layers.
+Building on this, we propose two complementary strategies—Recovery Testing (RT) and Logit-layer Testing (LT)—to empirically measure these violations and expose internal disruptions caused by adversaries.
+Evaluated on CIFAR-10, CIFAR-100, and ImageNet under both standard and adaptive threat models, our method achieves state-of-the-art detection performance with negligible computational overhead.
+Furthermore, our system-level analysis provides a practical method for selecting a detection threshold with a formal lower-bound guarantee on accuracy.
+The code is available here: https://github.com/c0510gy/AFLS-AED.
 Keywords Adversarial detection $\cdot$ Deep learning $\cdot$ Layer inconsistency $\cdot$ Robust defense $\cdot$ Adaptive attacks
@@
 Existing defensive strategies generally fall into three categories: adversarial training, input purification, and AE detection [15, 1].
-Adversarial training, re-trains or fine-tunes a model using adversarially perturbed samples [34, 20], despite its robustness [10, 36], it incurs high computational costs and may compromise clean-data accuracy.
+Adversarial training, re-trains or fine-tunes a model using adversarially perturbed samples to have Adversarially Trained Classifier (ATC) [35, 22], despite its robustness [11, 37], it incurs high computational costs and may compromise clean-data accuracy.
 Input purification methods [26, 32] attempt to remove adversarial noise through preprocessing (e.g., denoising), yet often fail against adaptive attacks [8].
-A more flexible alternative is AE detection, which avoids retraining or significantly modifying the model by rejecting suspicious inputs at test time [35, 40].
-Adversarial detection methods can be broadly categorized into two types: those that compare inputs to a reference set or distribution of known patterns (e.g., Deep k-Nearest Neighbors, DkNN [27]; Latent Neighborhood Graph, LNG [1]), and those that analyze invariants in the learned representations [15, 6].
-Reference-based detectors often require storing adversarial examples or constructing complex neighbor graphs, which can be computationally intensive and prone to overfitting specific attack types.
-To address these limitations, recent work [39] proposed detecting consistency between augmented inputs using pre-trained Self-Supervised Learning (SSL) models.
-Despite avoiding both the need for complex data structures and reliance on adversarial examples, it introduces several practical challenges.
-Notably, it incurs significant overhead due to the use of large pre-trained SSL models–often significantly larger (e.g., $\times2$ larger params) than the target classifier itself–and assumes the availability of high-quality SSL models for each domain, which may not always be feasible.
-Furthermore, their method relies solely on SSL models pre-trained on the same datasets used for evaluation.
+A more flexible alternative is AE detection, which avoids the costly retraining process of ATCs by instead rejecting suspicious inputs at test time [36, 41].
+Detection-based defenses offer distinct advantages, including lower implementation costs, as they do not require exhaustive adversarial training with adversarial data, and tunable robustness, where sensitivity can be adjusted to meet application-specific accuracy–robustness trade-offs.
+Furthermore, their plug-and-play nature allows them to be integrated with existing models, including ATCs, to further enhance system-level robustness.
+By architecturally separating the detection mechanism from the classifier, this approach introduces an additional layer of security at the system level.
+AE detection techniques are often categorized into two types: those that compare inputs to a reference set (e.g., Deep k-Nearest Neighbors, DkNN [29]; Latent Neighborhood Graph, LNG[1]), and those that analyze invariants in the learned representations [17, 7].
+Despite these benefits, prominent detection paradigms suffer from practical limitations.
+Reference-based detectors often require storing adversarial examples or constructing complex neighbor graphs, which is computationally intensive.
+To address this, recent work [40] has proposed detecting consistency between augmented inputs using pre-trained Self-Supervised Learning (SSL) models.
+However, this introduces significant overhead from large external models—which can be over 22 times larger than the target classifier—and assumes the availability of high-quality, domain-specific SSL models.
 To eliminate the overhead of external models, complex data structures, or heavy augmentations, we develop a self-contained, layer-wise detector that scrutinizes the network’s own representations.
 This raises a central question: How do adversarial perturbations propagate through a deep neural network?
-Prior work shows that such perturbations often leave early-layer features largely intact while inducing sharp deviations in deeper layers [15].
-Rather than contrasting these deviations with external reference sets or heavy augmented features, we ask our core research question: Do manifold-induced feature jumps themselves expose adversarial inputs?
-We formalize this intuition in the A Few Large Shifts Assumption, which posits that adversarial perturbations push representations onto fragile submanifolds at only a few critical layers, producing large, localized shifts.
-As illustrated in Figure 1, these sparse disruptions leave a detectable footprint across successive layers.
-Guided by this assumption, we introduce a lightweight detection framework grounded in layer-wise inconsistency.
-Two complementary probes—Recovery Testing (RT) for intermediate activations and Logit-layer Testing (LT) for the output layer—are calibrated solely on benign data.
-We then fuse their signals through Recovery-and-Logit Testing (RLT), aligning RT and LT scores via quantile normalization to robustly flag a broad spectrum of attacks, including those that evade one probe in isolation.
+Prior work suggests that such perturbations often leave early-layer features largely intact while inducing sharp deviations in deeper layers.
+Motivated by these findings, our own empirical analysis confirms this behavior (subsection 3.4).
+As visualized in Figure 3, we observe that while benign inputs produce relatively flat error profiles across layers, many common adversarial attacks induce large, localized shifts at only a few critical points in the network.
+Rather than contrasting these deviations with external references, we ask a different core question: Do these internal feature jumps themselves expose adversarial inputs?
+We formalize this intuition in the A Few Large Shifts Assumption, which posits that adversarial perturbations produce large, localized shifts between a few critical layers.
+To quantify the observed localized shifts, we introduce the concept of layer-wise Lipschitz continuity, defining these shifts as localized violations of layer-wise Lipschitz continuity.
+Guided by this principle, our framework is the first designed to empirically measure these violations using a detector calibrated only on benign data.
+As illustrated in Figure 1, these sparse disruptions leave a detectable footprint across successive layers, which we capture using a lightweight detection framework grounded in layer-wise inconsistency with two complementary probes: Recovery Testing (RT) and Logit-layer Testing (LT).
+We then fuse their signals through Recovery-and-Logit Testing (RLT), aligning their scores via quantile normalization to robustly flag a broad spectrum of attacks, including those that evade one probe in isolation.
 The distinguishing features and requirements of our method are summarized in Table 1, where we compare RT and LT against several representative baselines.
-As shown, our method is self-sufficient (requires no adversarial examples), model-local (requires no external SSL), and low-overhead (no kNN graphs or excessive augmentation), making it suitable for scalable deployment.
+As shown, our method is self-sufficient (requires no adversarial examples), model-local (requires no external SSL), and low-overhead (no kNN graphs or excessive augmentation).
+Furthermore, while prior AE detection works lack the thorough system-level analysis essential for practical deployment, our work addresses this gap by introducing a formal method to select a detection threshold with a guaranteed lower bound on overall system accuracy, as detailed in Appendix H.
+This enhances the practical reliability of our detector, making it highly suitable for scalable deployment.
 Our key contributions are as follows:
-  • We propose a novel layer-wise adversarial detection paradigm that exploits partial consistency in internal representations of adversarial examples by focusing on feature transformations.
-Unlike prior methods, our approach avoids significant overhead by eliminating the need for heavy augmentations, complex structures to maintain reference sets, external large pre-trained models, such as SSL, or large ensemble models.
-  • We introduce two practical testing strategies—Recovery Testing and Logit-layer Testing—which identify inconsistencies in intermediate and final-layer representations, respectively, using only the target model’s own feature space.
-  • We evaluate our framework on standard benchmarks (CIFAR-10, CIFAR-100, and ImageNet), demonstrating strong detection performance, low computational overhead, and negligible impact on clean inputs, making it practical for real-world deployment.
-By focusing on unique internal feature transformations over reference-based comparisons or heavy augmentations, our method avoids extensive retraining and specialized pipelines while delivering a more efficient, robust detector.
-Next, we review related work, detail our layer-consistency analysis, and present experiments across diverse adversarial threat models.
+  • We propose a novel adversarial detection paradigm that exploits partial consistency in internal feature transformations of adversarial examples.
+Unlike prior methods that rely on computationally expensive external reference sets or large auxiliary models, our approach eliminates this overhead by measuring inconsistencies across a model’s own layers for a single input.
+  • We are the first to formalize these internal disruptions as violations of layer-wise Lipschitz continuity.
+We introduce two complementary probes, Recovery Testing (RT) and Logit-layer Testing (LT), as practical and efficient methods to empirically quantify these violations using only benign data.
+  • We conduct a comprehensive evaluation on standard benchmarks, demonstrating state-of-the-art performance against a wide range of threats, including strong, end-to-end adaptive attacks and generalization to various architectures.
+Crucially, our system-level analysis yields a formal method for selecting a detection threshold with a guaranteed lower bound on system accuracy, enhancing the method’s practical reliability.
 Table 1: Comparison of Requirements / Properties Across Methods.
@@
 A   indicates that the method satisfies the criterion, while a   indicates it does not.
-Ours Baseline Methods Criterion RT LT BEYOND [39] LID [23] Mao [24] Hu [14] DkNN [27] kNN-Def.
-[9] Self-Sufficient Data?
+Ours Baseline Methods Criterion RT LT BEYOND LID Mao Hu DkNN kNN-Def.
+Self-Sufficient Data?
 (No Adversarial Data Needed?) Standalone Model?
@@
 (No Additional Training Needed?) /
-2 Related Work and Background
+Figure 2: Overview of our layer-wise adversarial detection framework.
+(Left) Recovery Testing (RT) trains regressors $R^{(L\rightarrow k)}$ to reconstruct intermediate features $z_{k}$ from the final embedding $z_{L}$; detection is based on the entropy-weighted reconstruction error $\|z_{k}-R^{(L\rightarrow k)}(z_{L})\|_{2}^{2}$.
+(Right) Logit-layer Testing (LT) applies perturbations $W^{(g)}$ to compute feature and logit discrepancies, $\Delta z^{(g)}=\|z_{i}-z_{i}^{(g)}\|_{2}^{2}$ and $\Delta\ell^{(g)}=\|o(\sigma(\ell))-\sigma(\ell^{(g)})\|_{2}^{2}$, which are combined as $\frac{H(\sigma(\ell))\Delta\ell^{(g)}}{\Delta z^{(g)}}$.
+Final RT and LT scores are averaged over $g$.
+(Bottom) Recovery and Logit Testing (RLT) integrates RT and LT scores using a linear combination after quantile normalization.
+2 Methodology
+In this section, we introduce our layer-wise adversarial example detection framework (Figure 2), comprising three measures: RT, LT, and their fusion RLT.
+RT captures inconsistencies in intermediate-layer embeddings, while LT measures logit instability under augmented inputs.
+At test time, pick a detection measure $m\in\{\mathrm{RT},\mathrm{LT},\mathrm{RLT}\}$ and threshold $\tau_{m}$.
+We declare $x$ adversarial iff $m(x)>\tau_{m}$, i.e., $\hat{a}(x)\;=\;\mathbb{I}\{\,m(x)>\tau_{m}\,\}\in\{0,1\}.$ Details of the selection of each threshold $\tau_{m}$ are delineated in Appendix H.
+The algorithm for calculating each measure is shown in Algorithm 1.
+Theoretical assumptions and proofs of our method are shown in Appendix B and Appendix C, respectively.
+2.1 Definition and Notations
+Given a target DNN $f(\cdot)=f_{logit}\circ f_{L}\circ f_{L-1}\circ\cdots\circ f_{1}(\cdot)$ with $L$ intermediate layers $f_{i}$ and one logit layer $f_{logit}$ at the end, we compute intermediate representations $z_{i}(x_{j})=f_{i}\circ\cdots\circ f_{1}(x_{j})\in\mathbb{R}^{D_{i}}$ for all $i\in\{1,2,\cdots,L\}$, where $D_{i}$ denotes the dimensionality of the embedding space and $x_{j}$ denotes an input sample.
+We assume the logit layer takes the last intermediate embedding $z_{L}$ to produce the final logit output $\ell=f_{logit}(z_{L})\in\mathbb{R}^{C}$, where $C$ is the number of output classes.
+We denote the softmax function as $\sigma(x)=\frac{\exp{x}}{\mathbf{1}^{\top}\exp{x}}$, the Shannon entropy as $\mathcal{H}(p)=-\sum p_{i}\log p_{i}$, and the one-hot vector of the ground-truth label $y$ as $\mathbf{o}_{y}\in\{0,1\}^{C}$.
+Throughout, all auxiliary modules are trained using only benign samples $\mathcal{D}_{norm}=\{x_{1},x_{2},\cdots,x_{N}\}$ without using labels.
+2.2 Core Assumption
+We begin by formalizing the central structural assumption underlying our detection framework:
+Assumption 1 (A Few Large Shifts).
+Let $x^{\mathrm{adv}}=x+\delta$ be an adversarial input, and let $z_{k}(x)$ denote the output of $k$-th layer $f_{k}$.
+We assume there exists a small subset of layers $\mathcal{T}\subseteq\{f_{1},\dots,f_{L},f_{logit}\}$ such that for any $f_{k}\in\mathcal{T}$,
+$\frac{1}{D_{k+1}}\|z_{k+1}(x^{\mathrm{adv}})-z_{k+1}(x)\|_{2}^{2}\gg\frac{1}{D_{k}}\|z_{k}(x^{\mathrm{adv}})-z_{k}(x)\|_{2}^{2}.$
+For all other layers $f_{j}\notin\mathcal{T}$, the shift between representations remains small:
+$\frac{1}{D_{j+1}}\|z_{j+1}(x^{\mathrm{adv}})-z_{j+1}(x)\|_{2}^{2}\approx\frac{1}{D_{j}}\|z_{j}(x^{\mathrm{adv}})-z_{j}(x)\|_{2}^{2}.$
+This assumption formalizes the intuition that adversarial perturbations cause disproportionately large changes in representation between specific adjacent layers while most other transitions remain relatively stable—a phenomenon that can be quantified based on Lipschitz continuity, which we term layer-wise Lipschitz continuity (See Appendix B.1).
+Specifically, the condition $\|\mathbf{z}_{k+1}(x^{\prime})-\mathbf{z}_{k+1}(x)\|_{2}\gg\|\mathbf{z}_{k}(x^{\prime})-\mathbf{z}_{k}(x)\|_{2}$ implies a large local Lipschitz constant for the transformation between layers, indicating a region of high instability.
+The following sections introduce our proposed testing measures, which are designed to empirically identify and quantify these localized violations of layer-wise Lipschitz continuity.
+2.3 Recovery Testing Measure
+For each hidden layer $k\in\{\,k_{RT},\,k_{RT}+1,\,\ldots,\,L-1\}$, we train an approximate inverse function: $R^{(L\rightarrow k)}:\mathbb{R}^{D_{L}}\longrightarrow\mathbb{R}^{D_{k}}$ to reconstruct $z_{k}$ from $z_{L}$ with hyperparameter $k_{RT}$.
+Each $R^{(L\rightarrow k)}$ is implemented as a lightweight MLP with 3–4 layers and trained by a mean squared error loss:
+$\mathcal{L}_{RT}=\frac{1}{N}\sum_{n=1}^{N}\sum_{k=k_{RT}}^{L-1}\|\,z_{k}(x_{n})-R^{(L\rightarrow k)}\bigl(z_{L}(x_{n})\bigr)\|_{2}^{2}.$ (1)
+At test time, we define the normalized squared error for each layer $k$ as
+$e_{k}(x)$ $=\|\,z_{k}(x)-R^{(L\rightarrow k)}(z_{L}(x))\|_{2}^{2},(k=1,\ldots,L-1.)$ (2)
+Collecting the vector $e(x)=(e_{1}(x),\ldots,e_{L-1}(x))$, we normalize it with softmax to obtain a distribution $\sigma(e(x))$ across layers.
+The non‐uniformity of this distribution is captured using inverse entropy $\log(L-1)-\mathcal{H}(\sigma(e(x)))$.
+The final score is computed by taking the average of the raw reconstruction errors and weighting it by this information‐based term:
+$RT(x)$ $=\bigl(\log(L-1)-\mathcal{H}(\sigma(e(x)))\bigr)\log\Bigl(\frac{1}{L-k_{RT}}\sum_{k=k_{RT}}^{L-1}e_{k}(x)\Bigr).$ (3)
+We apply a logarithm to the average of raw errors for numerical stability.
+Under our assumption, adversarial samples that target specific layers $\mathcal{T}\subseteq\{f_{1},\ldots,f_{L-1}\}$ will yield higher $RT(x)$ values due to large reconstruction errors (see Theorem 1) and a sharply peaked error distribution.
+Intuitively, the first term of Equation 3 measures how deterministic the error distribution is across layers, while the second term captures the magnitude of a few large shifts.
+2.4 Logit-layer Testing Measure
+Although RT effectively detects perturbations in intermediate layers $f_{i}$, it cannot be applied to the final logit layer $f_{\text{logit}}$, since there is no subsequent representation from which to reconstruct the logits.
+To address this—and motivated by the intuition that adversarial examples aim to induce misclassification—we introduce Logit‐layer Testing (LT): we quantify uncertainty at $f_{\text{logit}}$, relative to changes in intermediate features, using data‐driven, low‐cost input augmentations.
+Let $\{W^{(g)}\}_{g=1}^{G}$ be a small set (e.g., $1\leq G\leq 6$) of image transformation matrices.
+Each matrix $W^{(g)}$ is initialized as an identity transformation and is then fine-tuned on benign data using gradient descent.
+The LT score itself, as defined in Equation 4, serves as the loss function for this process, encouraging the transformations to find augmentations that preserve logit stability for benign inputs, i.e., $\sigma(f(x))\approx\sigma(f(W^{(g)}x))$, while still perturbing intermediate features.
+This establishes a consistent baseline against which adversarial instability can be measured.
+Given a test input, for each augmentation $g$, we measure two types of inconsistencies between the original and augmented inputs: feature-space drift $\Delta z^{(g)}(x|W^{(g)})$ and change in logit decidedness $\Delta\ell^{(g)}(x|W^{(g)})$.
+The feature-space drift is measured by averaging the $L_{2}$ distances across all intermediate outputs: $\Delta z^{(g)}(x|W^{(g)})=\frac{1}{L-k_{LT}+1}\sum_{i=k_{LT}}^{L}\|z_{i}(x)-z_{i}(W^{(g)}x)\|_{2}^{2}$ with hyperparameter $k_{LT}$.
+The logit decidedness is measured by computing the $L_{2}$ distance between the predicted one-hot vector and the softmax output: $\Delta\ell^{(g)}(x|W^{(g)})=\|\mathbf{o}_{\hat{y}}-\sigma(f(W^{(g)}x))\|_{2}^{2}$, where $\hat{y}=\arg\max_{c}\sigma_{c}(x)$.
+We then combine the two quantities into $\frac{\Delta\ell^{(g)}(x|W^{(g)})}{\Delta z^{(g)}(x|W^{(g)})}$, weight it by the entropy of the logit score vector $\mathcal{H}(\sigma(\ell(x)))$ to emphasize inputs closer to the decision boundary [13], and finally average over all augmentations as follows:
+$LT(x)=\frac{1}{G}\sum_{g=1}^{G}{\log{\mathcal{H}(\sigma(\ell(x)))\Delta\ell^{(g)}(x|W^{(g)})}-\log{\Delta z^{(g)}(x|W^{(g)})}}.$ (4)
+As before, the logarithm ensures numerical stability.
+Intuitively, under a successful adversarial attack the logit shift $\Delta\ell^{(g)}$ becomes disproportionately large compared to the accumulated feature drift $\Delta z^{(g)}$, which is precisely what LT is designed to detect (Theorem 3).
+To fine-tune the transformation matrices $\{W^{(g)}\}$, we directly use $LT(x)$ as the training loss:
+$\mathcal{L}_{LT}=\frac{1}{N}\sum_{n=1}^{N}LT(x_{n}).$ (5)
+2.5 Recovery and Logit Testing Combined Measure
+To capture inconsistencies in both intermediate and logit layers, we introduce a combined score called Recovery and Logit Testing.
+This score integrates RT and LT while correcting for differences in their statistical distributions using quantile normalization.
+Specifically, we transform each score based on its empirical cumulative distribution estimated from the benign training set, and map it to the standard normal distribution.
+Let $\hat{\mathcal{F}}_{RT}$ and $\hat{\mathcal{F}}_{LT}$ be the empirical cumulative distribution functions (CDFs) of RT and LT scores computed over the benign data.
+Let $\Phi^{-1}$ denote the quantile function (inverse CDF) of the standard normal distribution.
+At test time, each score is transformed as follows:
+$RT_{norm}(x)=\Phi^{-1}\left(\hat{\mathcal{F}}_{RT}(RT(x))\right),\quad LT_{norm}(x)=\Phi^{-1}\left(\hat{\mathcal{F}}_{LT}(LT(x))\right).$
+The final RLT score is computed by summing the squared normalized values:
+$RLT(x)=(RT_{norm}(x))^{2}+(LT_{norm}(x))^{2}.$ (6)
+This quantile-based transformation aligns both RT and LT scores to a common standard normal distribution, ensuring that the final combined score reflects significant deviations under either test (Theorem 4), independent of their original scales.
+This enhances robustness to varied attack types and scoring dynamics.
+3 Experiments
+Table 2: The AUC of Different Adversarial Detection Approaches on CIFAR-10.
+The results are the mean and standard deviation of 5 runs.
+Our methods are included for comparison.
+Classifier: ResNet110, FGSM: $\epsilon=0.05$, PGD: $\epsilon=0.02$.
+Note that our methods and BEYOND need no AE for training, leading to the same value on both seen and unseen settings.
+The bolded values are the best performance, and the underlined italicized values are the second-best performance.
+Method Unseen (Attacks in training are excluded from tests) Seen (Attacks in training are included in tests)
+FGSM PGD AutoAttack Square FGSM PGD CW AutoAttack Square
+DkNN [29] 61.55 ±0.023 51.22 ±0.026 52.12 ±0.023 59.46 ±0.022 61.55 ±0.023 51.22 ±0.026 61.52 ±0.028 52.12 ±0.023 59.46 ±0.022
+kNN [10] 61.83 ±0.018 54.52 ±0.022 52.67 ±0.022 73.39 ±0.020 61.83 ±0.018 54.52 ±0.022 62.23 ±0.019 52.67 ±0.022 73.39 ±0.020
+LID [25] 71.08 ±0.024 61.33 ±0.025 55.56 ±0.021 66.18 ±0.025 73.61 ±0.020 67.98 ±0.020 55.68 ±0.021 56.33 ±0.024 85.94 ±0.018
+Hu [16] 84.51 ±0.025 58.59 ±0.028 53.55 ±0.029 95.82 ±0.020 84.51 ±0.025 58.59 ±0.028 91.02 ±0.022 53.55 ±0.029 95.82 ±0.020
+Mao [26] 95.33 ±0.012 82.61 ±0.016 81.95 ±0.020 85.76 ±0.019 95.33 ±0.012 82.61 ±0.016 83.10 ±0.018 81.95 ±0.020 85.76 ±0.019
+LNG [1] 98.51 63.14 58.47 94.71 99.88 91.39 89.74 84.03 98.82
+BEYOND [40] 98.89 ±0.013 99.28 ±0.020 99.16 ±0.021 99.27 ±0.016 98.89 ±0.013 99.28 ±0.020 99.20 ±0.008 99.16 ±0.021 99.27 ±0.016
+Our Approaches
+RT 99.93 ±0.005 96.89 ±0.071 99.99 ±0.000 85.38 ±0.344 99.93 ±0.005 96.89 ±0.071 99.99 ±0.002 99.99 ±0.000 85.38 ±0.344
+LT 97.50 ±0.038 98.61 ±0.042 99.60 ±0.018 97.47 ±0.036 97.50 ±0.038 98.61 ±0.042 97.08 ±0.027 99.60 ±0.018 97.47 ±0.036
+RLT 99.85 ±0.005 99.37 ±0.011 99.99 ±0.000 95.95 ±0.102 99.85 ±0.005 99.37 ±0.011 99.91 ±0.004 99.99 ±0.000 95.95 ±0.102
+Table 3: The AUC of Different Adversarial Detection Approaches on ImageNet.
+To align with baselines, classifier: DenseNet121, FGSM: $\epsilon=0.05$, PGD: $\epsilon=0.02$.
+Due to memory and resource constraints, baseline methods are not evaluated against AutoAttack on ImageNet.
+Method Unseen Seen
+FGSM PGD FGSM PGD CW
+DkNN [29] 89.16 ±0.038 78.00 ±0.041 89.16 ±0.038 78.00 ±0.041 68.91 ±0.044
+kNN [10] 51.63 ±0.04 51.14 ±0.039 51.63 ±0.04 51.14 ±0.039 50.73 ±0.04
+LID [25] 90.32 ±0.046 52.56 ±0.038 99.24 ±0.043 98.09 ±0.042 58.83 ±0.041
+Hu [16] 72.56 ±0.037 86.00 ±0.042 72.56 ±0.037 86.00 ±0.042 80.79 ±0.044
+LNG [1] 96.85 89.61 99.53 98.42 86.05
+BEYOND [40] 97.59 ±0.04 96.26 ±0.045 97.59 ±0.04 96.26 ±0.045 95.46 ±0.047
+Our Approaches
+RT 94.31 ±0.457 99.99 ±0.000 94.31 ±0.457 99.99 ±0.000 92.18 ±0.135
+LT 96.18 ±0.028 97.89 ±0.021 96.18 ±0.028 97.89 ±0.021 94.06 ±0.215
+RLT 97.60 ±0.048 99.99 ±0.000 97.60 ±0.048 99.99 ±0.000 91.19 ±0.022
+Table 4: RA(%) under Orthogonal-PGD Adaptive Attack using CIFAR-10 and ResNet110.
+Defense $\bm{L_{\infty}=0.01}$ $\bm{L_{\infty}=8/255}$
+RA@FPR5% RA@FPR50% RA@FPR5% RA@FPR50%
+Ours (RLT) 75.40 99.58 33.70 80.77
+BEYOND [40] 88.38 98.81 13.80 48.20
+Trapdoor [31] 0.00 7.00 0.00 8.00
+DLA [33] 62.60 83.70 0.00 28.20
+SID [34] 6.90 23.40 0.00 1.60
+SPAM [21] 1.20 46.00 0.00 38.00
+Table 5: RA (%) scores of end-to-end attack with PGD under perturbation $\epsilon=8/255$ (ResNet110, CIFAR10). The results are the mean and standard deviation of 5 runs.
+$\bm{\lambda}$ @FPR5% @FPR10% @FPR15% @FPR20% @FPR25% @FPR30% @FPR35% @FPR40% @FPR45% @FPR50%
+1.00 $20.72_{\pm 8.04}$ $25.96_{\pm 6.88}$ $32.05_{\pm 5.72}$ $38.23_{\pm 5.11}$ $44.48_{\pm 4.52}$ $0.15_{\pm 4.13} 55$$5.15_{\pm 3.56}$ $60.09_{\pm 3.10}$ $64.54_{\pm 2.42}$ $68.53_{\pm 2.21}$
+0.50 $22.64_{\pm 4.46}$ $30.45_{\pm 3.25}$ $37.96_{\pm 2.08}$ $44.78_{\pm 1.81}$ $51.19_{\pm 1.30}$ $56.41_{\pm 0.98}$ $61.21_{\pm 0.96}$ $65.44_{\pm 0.96}$ $69.35_{\pm 0.83}$ $73.00_{\pm 0.92}$
+0.25 $23.39_{\pm 1.76}$ $32.92_{\pm 0.49}$ $40.88_{\pm 1.04}$ $48.04_{\pm 1.13}$ $54.32_{\pm 1.58}$ $59.37_{\pm 1.44}$ $63.90_{\pm 1.82}$ $67.98_{\pm 1.82}$ $71.51_{\pm 1.89}$ $74.83_{\pm 1.85}$
+Mao [26] 18.97
+BEYOND [40] 19.45
+Table 6: RA (%) scores of end-to-end attack with PGD under perturbation $\epsilon=8/255$ (ResNet110, CIFAR10) when using an ATC. The results are the mean and standard deviation of 5 runs.
+$\bm{\lambda}$ @FPR5% @FPR10% @FPR15% @FPR20% @FPR25% @FPR30% @FPR35% @FPR40% @FPR45% @FPR50%
+1.00 $93.32_{\pm 0.13}$ $93.38_{\pm 0.14}$ $93.41_{\pm 0.11}$ $93.43_{\pm 0.09}$ $93.97_{\pm 0.12}$ $94.49_{\pm 0.15}$ $94.99_{\pm 0.14}$ $95.50_{\pm 0.13}$ $95.93_{\pm 0.14}$ $96.32_{\pm 0.11}$
+0.50 $93.53_{\pm 0.15}$ $94.17_{\pm 0.12}$ $94.27_{\pm 0.11}$ $94.33_{\pm 0.09}$ $94.86_{\pm 0.13}$ $95.41_{\pm 0.09}$ $95.88_{\pm 0.09}$ $96.27_{\pm 0.11}$ $96.65_{\pm 0.14}$ $97.00_{\pm 0.15}$
+0.25 $93.57_{\pm 0.14}$ $94.46_{\pm 0.16}$ $94.60_{\pm 0.13}$ $94.85_{\pm 0.07}$ $95.40_{\pm 0.05}$ $95.87_{\pm 0.08}$ $96.32_{\pm 0.14}$ $96.71_{\pm 0.12}$ $97.04_{\pm 0.12}$ $97.30_{\pm 0.14}$
+Mao [26] 75.09
+BEYOND [40] 93.20
+Table 7: Evaluation of RA (%) on end-to-end gradient-free SimBA attack with 1000 steps under perturbation $\epsilon=8/255$ (ResNet110, CIFAR10).
+Defense RA@FPR5% RA@FPR50%
+None 5.54
+Ours (RLT) 97.62 97.85
+Figure 3: Empirical validation of the A Few Large Shifts Assumption using CIFAR-10 and ResNet-110. We plot layer-wise error distributions $\sigma(e_{k})$ under different attack methods.
+Table 8: Comparison of implementation costs in terms of FLOPs, parameters, and model size overhead when the target model is ResNet110 with the CIFAR-10 dataset.
+[…]
+Mao [26] 5.25 38.12 22.02
+LNG [1] 0.286 8.33 4.81
+BEYOND [40] 0.715 20.62 11.91
+Ours (RLT) 0.491 2.59 1.49
+3.1 Experimental Setup
+Datasets.
+We evaluate our method on both CIFAR-10 and ImageNet datasets.
+CIFAR-10 consists of 60,00060{,}000 $32\times 32$ images across 10 classes, with standard training (50k) and test (10k) splits.
+For ImageNet, we use the official training and validation sets, resizing all images to $256\times 256$ and applying standard normalization.
+CIFAR-10 is used to validate the generality and scalability of our detection approach in a lower-resolution setting, while ImageNet serves as a large-scale benchmark.
+Backbone Models.
+For CIFAR-10, we use a ResNet-110 pretrained classifier.
+For ImageNet, we employ a pretrained DenseNet-121.
+In both cases, we extract intermediate features from multiple layers without fine-tuning the backbone, ensuring a consistent foundation for detection.
+Threat Models.
+We evaluate our detection framework under two standard adversarial settings: Limited Knowledge and Perfect Knowledge, following the protocol of previous works [3, 40].
+In the Limited Knowledge setting, the adversary has full access to the target classifier but is unaware of the detection mechanism, which remains confidential.
+In contrast, the Perfect Knowledge (adaptive attack) setting assumes that the adversary has full knowledge of both the classifier and the detection strategy, enabling it to craft attacks specifically to evade detection.
+Attack Methods.
+We evaluate robustness and detection performance under a diverse suite of white-box adversarial attacks, including Fast Gradient Sign Method (FGSM), Projected Gradient Descent (PGD), and Carlini–Wagner (CW) attacks.
+We also apply AutoAttack (standard version) and Square Attack [2], both of which are parameter-free (black-box) and ensemble-based.
+To assess robustness under this stronger threat model, we incorporate the state-of-the-art adaptive attack, Orthogonal-PGD [5], which optimizes adversarial perturbations not only to fool the classifier but also to minimize detection signals.
+This provides a rigorous evaluation of the proposed methods under any conditions.
+Implementation Details.
+All models are implemented in PyTorch and trained on NVIDIA RTX 6000 Ada.
+We use a batch size of 32, AdamW optimizer with learning rate $1\times 10^{-4}$ and weight decay of $0.01$, training the recover modules and augmentation matrices for 50 epochs.
+3.2 Detection Performance under Standard Attacks
+We evaluate the effectiveness of our proposed detection scores, namely RT, LT, and RLT, on CIFAR-10 and ImageNet under standard adversarial threat models, including FGSM, PGD, CW, AutoAttack, and Square Attack.
+We benchmark against established baselines such as LID [25], DkNN [29], LNG [1], and the recent SSL-based BEYOND [40].
+The Area Under the Receiver Operating Characteristic Curve (RoC-AUC) results, which measure detection performance on various thresholds, are reported in Table 2 for CIFAR-10 and Table 3 for ImageNet.
+On CIFAR-10, RT achieves exceptionally strong performance across attacks that disrupt internal representations (e.g., FGSM, CW, AutoAttack), aligning with its role in capturing large, localized deviations in intermediate layers.
+LT performs particularly well against attacks like PGD and Square, which introduce minimal internal distortion but induce instability at the output layer.
+Notably, the fused RLT score achieves either the best or second-best AUC across most of the attacks, validating the complementary nature of RT and LT.
+Despite relying solely on the internal signals of the target classifier and requiring no adversarial examples or external models, RLT matches or outperforms state-of-the-art methods, including BEYOND, which uses large pre-trained SSL representations.
+On ImageNet, we observe similar trends.
+RT continues to perform well under strong perturbations such as PGD, while LT maintains robust accuracy under less structured noise like FGSM.
+The combined RLT score again leads to superior or competitive detection performance.
+Unlike several baselines that degrade substantially in the large-scale setting due to increased model capacity or overfitting to specific attacks, our method remains stable without requiring additional training resources or architecture-specific tuning.
+3.3 Detection Performance under Adaptive Attacks
+To avoid gradient obfuscation and guarantee end-to-end gradient flow through our detection framework, we apply the Backward Pass Differentiable Approximation (BPDA) [4] to components that may otherwise block gradients, such as the quantile module.
+We first evaluate robustness under the Orthogonal-PGD adaptive attack following BEYOND [40], where the adversary has full knowledge of both the classifier and our detection mechanism and explicitly optimizes to induce misclassification while suppressing detection.
+Table 4 reports robust accuracy (RA) at two $L_{\infty}$ budgets ($0.01$ and $8/255$) and operating points (FPR=5% and FPR=50%).
+While BEYOND benefits from additional SSL components, our method attains competitive or superior RA using only model-internal signals and benign calibration, without adversarial data or external encoders.
+This demonstrates that even when gradients are orthogonalized, the detection terms themselves impose inherently conflicting objectives, thereby constraining the attacker’s effectiveness.
+We further evaluate fully end-to-end PGD attacks on the fused RLT score to validate the analysis on objective confliction in Appendix B.5.2, using the untargeted loss $\mathcal{L}_{\mathrm{cls}}(x+\delta,y)+\lambda\cdot\mathrm{RLT}(x+\delta)$.
+Table 5 shows that, although RA decreases under this strong adaptive threat, our method consistently outperforms baselines’ best performance report at $\mathrm{FPR}{=}5\%$, such as Mao et al.
+and BEYOND across all tested $\lambda$ and FPRs.
+When combined with an ATC (Table 6), the system remains highly resilient, achieving over $93.3\%$ RA at $\mathrm{FPR}{=}5\%$, again surpassing the adaptive-attack results of Mao et al.
+and BEYOND at $\mathrm{FPR}{=}5\%$.
+These findings underscore that even under fully end-to-end optimization with varying attack strengths $\lambda$, robustness is maintained through gradient conflicts induced by our fused detection terms.
+To rule out gradient masking completely, we evaluate a query-only, gradient-free attack (SimBA) [14] with 1,000 queries.
+Table 7 shows that although SimBA reduces the undefended classifier’s RA to 5.54%, our RLT detector restores RA to 97.62% at 5% FPR.
+This strong performance against a computationally intensive black-box attack confirms that our defense does not rely on obfuscated gradients and remains robust beyond first-order threat models.
+3.4 Empirical Evaluation of the Proposed Assumption
+To validate our A Few Large Shifts Assumption, we measured layer-wise reconstruction error distributions $\sigma(e_{k})$ on a ResNet-110 model trained on CIFAR-10 under various attacks: FGSM, PGD, CW, AutoAttack, and Square Attack.
+As shown in Figure 3, benign inputs exhibit relatively flat error profiles across layers, while certain attacks as shown in Figure 3.(a), produce sharp peaks at specific layers—indicating large indicated as red arrows, localized shifts in internal representations.
+These peaked patterns align with high RT detection performance and support the assumption that adversarial perturbations affect only a subset of layers $\mathcal{T}\subseteq\{f_{1},f_{2},\cdots,f_{L}\}$.
+In contrast, attacks like PGD and Square, which show stronger LT detection, yield smoother error curves similar to benign inputs but with greater variance as depicted in Figure 3.(b).
+This suggests that these attacks primarily affect the logit layer, with minimal disruption to intermediate features $\mathcal{T}=\{f_{logit}\}$.
+These findings empirically confirm that adversarial manipulations often concentrate at a few layers and further justify the complementary roles of RT and LT in our detection framework.
+3.5 Implementation Costs
+We evaluate the computational efficiency of our RLT method relative to existing baselines under the same configuration as in Table 2.
+As shown in Table 8, RLT introduces minimal floating-point operations (FLOPs), parameters, and model size overhead, offering a lightweight, plug-in solution without retraining or external models.
+It is significantly more efficient than SSL-based (Mao et al.[26]) and graph-based (LNG[1]) detectors, making it well-suited for real-time and resource-constrained deployments.
+A detailed analysis of this detection overhead across various architectures is provided in Appendix G.
+4 Conclusion
+In this work, we introduced a novel framework for adversarial example detection that leverages layer-wise inconsistencies within deep neural networks.
+Motivated by our proposed A Few Large Shifts Assumption—which posits that adversarial perturbations cause large, localized violations of layer-wise Lipschitz continuity—we developed two complementary detection strategies: RT and LT.
+Extensive evaluations on CIFAR-10, CIFAR-100, and ImageNet demonstrate that our combined approach, RLT, achieves state-of-the-art detection performance against a wide range of threats, including strong adaptive attacks.
+Our method operates efficiently without relying on external models or extensive augmentation, and our system-level analysis provides a practical method for selecting a detection threshold with a formal lower-bound guarantee on accuracy, highlighting its suitability for real-world deployment.
+Limitations and Future Work.
+Despite these results, the framework’s effectiveness relies on an assumption grounded in empirical observations of current attacks and thus cannot guarantee defense against all possible future attacks.
+Key directions for future work therefore include deepening the theoretical understanding of layer-wise Lipschitz continuity to determine if the "A Few Large Shifts" principle is a fundamental property of adversarial examples, and developing more sophisticated methods for analyzing the challenging logit space.
+This exploration could lead to provably robust defenses.
+Acknowledgements
+This work was supported in part by the DARPA Young Faculty Award, the National Science Foundation (NSF) under Grants #2127780, #2319198, #2321840, #2312517, and #2235472, the Semiconductor Research Corporation (SRC), the Office of Naval Research through the Young Investigator Program Award, and Grants #N00014-21-1-2225 and #N00014-22-1-2067, Army Research Office Grant #W911NF2410360, and the U.S.
+Army Combat Capabilities Development Command (DEVCOM) Army Research Laboratory under Support Agreement No.
+USMA 21050.
+Additionally, support was provided by the Air Force Office of Scientific Research under Award #FA9550-22-1-0253, along with generous gifts from Xilinx and Cisco.
+The views expressed in this paper are those of the authors and do not reflect the official policy or position of the U.S.
+Military Academy, the U.S.
+Army, the U.S.
+Department of Defense, or the U.S.
+Government.
+Appendix A Related Work and Background
 Input Transformation-Based Detectors.
@@
 For example, BEYOND [40] detects adversarial inputs by measuring feature stability across augmentations in a large SSL encoder.
-These methods eliminate the need for adversarial examples and external graph structures but introduce high computational costs due to reliance on large pretrained models, which may be significantly larger than the target classifier itself (e.g., $\times2$ larger parameters).
+These methods eliminate the need for adversarial examples and external graph structures but introduce high computational costs due to reliance on large pretrained models, which may be significantly larger than the target classifier itself (e.g., $22\times$ larger parameters).
 Additionally, they often require access to domain-specific SSL models and retraining or fine-tuning of auxiliary heads.
 Our Approach.
-In contrast to the methods above, our framework proposes a lightweight, attack-agnostic detection mechanism that operates directly on the internal representations of a pre-trained target classifier.
-It builds on the assumption that adversarial perturbations typically cause large deviations at a few critical layers (A Few Large Shifts Assumption), Our method eliminates the need for adversarial examples, auxiliary models, and graph-based structures by relying solely on benign calibration data and a handful of lightweight augmentations (< 6).
-By emphasizing feature evolution across layers rather than comparing to external references, we realize a simple, highly efficient detector.
-Extensive evaluations on CIFAR-10, CIFAR-100, and ImageNet demonstrate strong detection performance with efficient and robust defense against both adaptive and white-box attacks.
-Figure 2: Overview of our layer-wise adversarial detection framework.
-(Left) Recovery Testing (RT) trains regressors $R^{(L\rightarrow k)}$ to reconstruct intermediate features $z_{k}$ from the final embedding $z_{L}$; detection is based on the entropy-weighted reconstruction error $\|z_{k}-R^{(L\rightarrow k)}(z_{L})\|_{2}^{2}$.
-(Right) Logit-layer Testing (LT) applies perturbations $W^{(g)}$ to compute feature and logit discrepancies, $\Delta z^{(g)}=\|z_{i}-z_{i}^{(g)}\|_{2}^{2}$ and $\Delta\ell^{(g)}=\|o(\sigma(\ell))-\sigma(\ell^{(g)})\|_{2}^{2}$, which are combined as $\frac{H(\sigma(\ell))\Delta\ell^{(g)}}{\Delta z^{(g)}}$.
-Final RT and LT scores are averaged over $g$.
-(Bottom) Recovery and Logit Testing (RLT) integrates RT and LT scores using a linear combination after quantile normalization.
-3 Methodology
-Algorithm 1 Layer-wise Adversarial Detection Measures via RT, LT, and RLT
-$f=f_{logit}\circ f_{L}\circ\dots\circ f_{1}$: Target network
-2:$\{R^{(L\rightarrow k)}\}_{k=k_{RT}}^{L-1}$: Trained inverse regressors
-3:$\{W^{(g)}\}_{g=1}^{G}$: Learned augmentation matrices
-4:^$\hat{\mathcal{F}}_{RT},\hat{\mathcal{F}}_{LT}$: Empirical CDFs of RT and LT (from benign data)
-5:$\Phi^{-1}$: Standard normal quantile function
-6:$x$: Test input
-7:
-8:function RT($x$)
-9:     for $k=k_{RT}$ to $L-1$ do
-10:         $e_{k}\leftarrow\|z_{k}(x)-R^{(L\rightarrow k)}(z_{L}(x))\|_{2}^{2}$
-11:     end for
-12:     $\bm{e}\leftarrow(e_{1},\dots,e_{L-1})$
-13:     return $(\log(L-1)-\mathcal{H}(\sigma(\bm{e})))\cdot\log\left(\frac{1}{L-1}\sum_{k}e_{k}\right)$
-14:end function
-15:
-16:function LT($x$)
-17:     for $g=1$ to $G$ do
-18:         $\Delta z^{(g)}\leftarrow\frac{1}{L-k_{LT}+1}\sum_{i=k_{LT}}^{L}\|z_{i}(x)-z_{i}(W^{(g)}x)\|_{2}^{2}$
-19:         $\hat{y}\leftarrow\arg\max\sigma(f(x))$
-20:         $\Delta\ell^{(g)}\leftarrow\|\mathbf{o}_{\hat{y}}-\sigma(f(W^{(g)}x))\|_{2}^{2}$
-21:         $s^{(g)}\leftarrow\log\left(\mathcal{H}(\sigma(f(x)))\cdot\Delta\ell^{(g)}\right)-\log\Delta z^{(g)}$
-22:     end for
-23:     return $\frac{1}{G}\sum_{g=1}^{G}s^{(g)}$
-24:end function
-25:
-26:function RLT($x$)
-27:     $r\leftarrow$ RT($x$),  $l\leftarrow$ LT($x$)
-28:     $r_{norm}\leftarrow\Phi^{-1}(\hat{\mathcal{F}}_{RT}(r))$
-29:     $l_{norm}\leftarrow\Phi^{-1}(\hat{\mathcal{F}}_{LT}(l))$
-30:     return $r_{norm}^{2}+l_{norm}^{2}$
-31:end function
-In this section, we introduce our layer-wise adversarial example detection framework (Figure 2), comprising three measures: RT, LT, and their fusion RLT.
-RT captures inconsistencies in intermediate-layer embeddings, while LT measures logit instability under augmented inputs.
-At test time, given measure $m\in\{\mathrm{RT},\mathrm{LT},\mathrm{RLT}\}\quad\text{and}\quad\tau_{m},$ we classify $\hat{a}(x)=\mathbf{1}^{\top}\bigl{(}f(x)>\tau_{f}\bigr{)}\in\{0,1\}$ as AE detection results.
-The algorithm for calculating each measure is shown in Algorithm 1.
-Theoretical assumptions and proofs of our method are shown in Appendix A and Appendix B, respectively.
-3.1 Definition and Notations
-Given a target DNN $f(\cdot)=f_{logit}\circ f_{L}\circ f_{L-1}\circ\cdots\circ f_{1}(\cdot)$ with $L$ intermediate layers $f_{i}$ and one logit layer $f_{logit}$ at the end, we compute intermediate representations $z_{i}(x_{j})=f_{i}\circ\cdots\circ f_{1}(x_{j})\in\mathbb{R}^{D_{i}}$ for all $i\in\{1,2,\cdots,L\}$, where $D_{i}$ denotes the dimensionality of the embedding space and $x_{j}$ denotes an input sample.
-We assume the logit layer takes the last intermediate embedding $z_{L}$ to produce the final logit output $\ell=f_{logit}(z_{L})\in\mathbb{R}^{C}$, where $C$ is the number of output classes.
-We denote the softmax function as $\sigma(x)=\frac{\exp{x}}{\mathbf{1}^{\top}\exp{x}}$, the Shannon entropy as $\mathcal{H}(p)=-\sum p_{i}\log p_{i}$, and the one-hot vector of the ground-truth label $y$ as $\mathbf{o}_{y}\in\{0,1\}^{C}$.
-Throughout, all auxiliary modules are trained using only benign samples $\mathcal{D}_{norm}=\{x_{1},x_{2},\cdots,x_{N}\}$ without using labels.
-3.2 Core Assumption
-We begin by formalizing the central structural assumption underlying our detection framework:
-Assumption 1 (A Few Large Shifts).
-Let $x^{\mathrm{adv}}=x+\delta$ be an adversarial input, and let $z_{k}(x)$ denote the output of $k$-th layer $f_{k}$.
-We assume there exists a small subset of layers $\mathcal{T}\subseteq\{f_{1},\dots,f_{L},f_{logit}\}$ such that for any $f_{k}\in\mathcal{T}$,
-$\frac{1}{D_{k+1}}\|z_{k+1}(x^{\mathrm{adv}})-z_{k+1}(x)\|_{2}^{2}\gg\frac{1}{D_{k}}\|z_{k}(x^{\mathrm{adv}})-z_{k}(x)\|_{2}^{2}.$
-For all other layers $f_{j}\notin\mathcal{T}$, the shift between representations remains small:
-$\frac{1}{D_{j+1}}\|z_{j+1}(x^{\mathrm{adv}})-z_{j+1}(x)\|_{2}^{2}\approx\frac{1}{D_{j}}\|z_{j}(x^{\mathrm{adv}})-z_{j}(x)\|_{2}^{2}.$
-This assumption formalizes the intuition that adversarial perturbations cause disproportionately large changes in representation between specific adjacent layers, while most other transitions remain relatively stable.
-These localized disruptions manifest as sharp deviations in feature space at only a few critical points in the network.
-In the following sections, we introduce our proposed testing measures designed to capture and quantify these irregular transitions.
-3.3 Recovery Testing Measure
-For each hidden layer $k\in\{\,k_{RT},\,k_{RT}+1,\,\ldots,\,L-1\}$, we train an approximate inverse function: $R^{(L\rightarrow k)}:\mathbb{R}^{D_{L}}\longrightarrow\mathbb{R}^{D_{k}}$ to reconstruct $z_{k}$ from $z_{L}$ with hyperparameter $k_{RT}$.
-Each $R^{(L\rightarrow k)}$ is implemented as a lightweight MLP with 3–4 layers and trained by a mean squared error loss:
-$\mathcal{L}_{RT}=\frac{1}{N}\sum_{n=1}^{N}\sum_{k=k_{RT}}^{L-1}\|\,z_{k}(x_{n})-R^{(L\rightarrow k)}\bigl{(}z_{L}(x_{n})\bigr{)}\|_{2}^{2}.$ (1)
-At test time, we define the normalized squared error for each layer $k$ as
-$e_{k}(x)$ $=\|\,z_{k}(x)-R^{(L\rightarrow k)}(z_{L}(x))\|_{2}^{2},(k=1,\ldots,L-1.)$ (2)
-Collecting the vector $e(x)=(e_{1}(x),\ldots,e_{L-1}(x))$, we normalize it with softmax to obtain a distribution $\sigma(e(x))$ across layers.
-The non-uniformity of this distribution is captured using inverse entropy $\log(L-1)-\mathcal{H}(\sigma(e(x)))$.
-The final score is computed by taking the average of the raw reconstruction errors and weighting it by this information-based term:
-$RT(x)$ $=\bigl{(}\log(L-1)-\mathcal{H}(\sigma(e(x)))\bigr{)}\log\Bigl{(}\frac{1}{L-k_{RT}}\sum_{k=k_{RT}}^{L-1}e_{k}(x)\Bigr{)}.$ (3)
-We apply a logarithm to the average of raw errors for numerical stability.
-Under our assumption, adversarial samples that target specific layers $\mathcal{T}\subseteq\{f_{1},\ldots,f_{L-1}\}$ will yield higher $RT(x)$ values due to large reconstruction errors (see Theorem 1) and a sharply peaked error distribution.
-Intuitively, the first term of Equation 3 measures how deterministic the error distribution is across layers, while the second term captures the magnitude of a few large shifts.
-3.4 Logit-layer Testing Measure
-Although RT effectively detects perturbations in intermediate layers $f_{i}$, it cannot be applied to the final logit layer $f_{\text{logit}}$, since there is no subsequent representation from which to reconstruct the logits.
-To address this—and motivated by the intuition that adversarial examples aim to induce misclassification—we introduce Logit-layer Testing (LT): we quantify uncertainty at $f_{\text{logit}}$, relative to changes in intermediate features, using data-driven, low-cost input augmentations.
-Let $\{W^{(g)}\}_{g=1}^{G}$ be a small set (e.g., $1\leq G\leq 6$) of image transformation matrices.
-Each $W^{(g)}$ is fine-tuned on benign data by minimizing the difference in prediction results, i.e., $\sigma(f(x))\approx\sigma(f(W^{(g)}x))$, while inducing larger perturbations in the intermediate layers.
-This encourages finding $W^{(g)}$ such that $f_{logit}$ remains robust in its prediction confidence, even when the intermediate layers $f_{i}$ are less robust for benign inputs.
-Given a test input, for each augmentation $g$, we measure two types of inconsistencies between the original and augmented inputs: feature-space drift $\Delta z^{(g)}(x|W^{(g)})$ and change in logit decidedness $\Delta\ell^{(g)}(x|W^{(g)})$.
-The feature-space drift is measured by averaging the $L_{2}$ distances across all intermediate outputs: $\Delta z^{(g)}(x|W^{(g)})=\frac{1}{L-k_{LT}+1}\sum_{i=k_{LT}}^{L}\|z_{i}(x)-z_{i}(W^{(g)}x)\|_{2}^{2}$ with hyperparameter $k_{LT}$.
-The logit decidedness is measured by computing the $L_{2}$ distance between the predicted one-hot vector and the softmax output: $\Delta\ell^{(g)}(x|W^{(g)})=\|\mathbf{o}_{\hat{y}}-\sigma(f(W^{(g)}x))\|_{2}^{2}$, where $\hat{y}=\arg\max_{c}\sigma_{c}(x)$.
-We then combine the two quantities into $\frac{\Delta\ell^{(g)}(x|W^{(g)})}{\Delta z^{(g)}(x|W^{(g)})}$, weight it by the entropy of the logit score vector $\mathcal{H}(\sigma(\ell(x)))$ to emphasize inputs closer to the decision boundary [12], and finally average over all augmentations as follows:
-$LT(x)=\frac{1}{G}\sum_{g=1}^{G}{\log{\mathcal{H}(\sigma(\ell(x)))\Delta\ell^{(g)}(x|W^{(g)})}-\log{\Delta z^{(g)}(x|W^{(g)})}}.$ (4)
-As before, the logarithm ensures numerical stability.
-Intuitively, under a successful adversarial attack the logit shift $\Delta\ell^{(g)}$ becomes disproportionately large compared to the accumulated feature drift $\Delta z^{(g)}$, which is precisely what LT is designed to detect (Theorem 3).
-To fine-tune the transformation matrices $\{W^{(g)}\}$, we directly use $LT(x)$ as the training loss:
-$\mathcal{L}_{LT}=\frac{1}{N}\sum_{n=1}^{N}LT(x_{n}).$ (5)
-3.5 Recovery and Logit Testing Combined Measure
-To capture inconsistencies in both intermediate and logit layers, we introduce a combined score called Recovery and Logit Testing.
-This score integrates RT and LT while correcting for differences in their statistical distributions using quantile normalization.
-Specifically, we transform each score based on its empirical cumulative distribution estimated from the benign training set, and map it to the standard normal distribution.
-Let ^$\hat{\mathcal{F}}_{RT}$ and ^$\hat{\mathcal{F}}_{LT}$ be the empirical cumulative distribution functions (CDFs) of RT and LT scores computed over the benign data.
-Let $\Phi^{-1}$ denote the quantile function (inverse CDF) of the standard normal distribution.
-At test time, each score is transformed as follows:
-$RT_{norm}(x)=\Phi^{-1}\left(\hat{\mathcal{F}}_{RT}(RT(x))\right),\quad LT_{norm}(x)=\Phi^{-1}\left(\hat{\mathcal{F}}_{LT}(LT(x))\right).$
-The final RLT score is computed by summing the squared normalized values:
-$RLT(x)=(RT_{norm}(x))^{2}+(LT_{norm}(x))^{2}.$ (6)
-This quantile-based transformation aligns both RT and LT scores to a common standard normal distribution, ensuring that the final combined score reflects significant deviations under either test (Theorem 4), independent of their original scales.
-This enhances robustness to varied attack types and scoring dynamics.
-4 Experiments
-To evaluate the effectiveness of our proposed methods, we conduct five main experiments: standard adversarial detection on CIFAR-10 and ImageNet (subsection 4.2); robustness evaluation under adaptive attacks (Orthogonal-PGD) (subsection 4.3); empirical validation of our core assumption on internal feature shifts (subsection 4.4); analysis of computational overhead for efficiency (subsection 4.2); and an ablation study to assess the contribution of each component (Appendix C).
-4.1 Experimental Setup
-Datasets.
-We evaluate our method on both CIFAR-10 and ImageNet datasets.
-CIFAR-10 consists of $60{,}000$ $32\times 32$ images across 10 classes, with standard training (50k) and test (10k) splits.
-For ImageNet, we use the official training and validation sets, resizing all images to $256\times 256$ and applying standard normalization.
-CIFAR-10 is used to validate the generality and scalability of our detection approach in a lower-resolution setting, while ImageNet serves as a large-scale benchmark.
-Backbone Models.
-For CIFAR-10, we use a ResNet-110 pretrained classifier.
-For ImageNet, we employ a pretrained DenseNet-121.
-In both cases, we extract intermediate features from multiple layers without fine-tuning the backbone, ensuring a consistent foundation for detection.
-Threat Models.
-We evaluate our detection framework under two standard adversarial settings: Limited Knowledge and Perfect Knowledge, following the protocol of previous works [3, 39].
-In the Limited Knowledge setting, the adversary has full access to the target classifier but is unaware of the detection mechanism, which remains confidential.
-In contrast, the Perfect Knowledge (adaptive attack) setting assumes that the adversary has full knowledge of both the classifier and the detection strategy, enabling it to craft attacks specifically to evade detection.
-Attack Methods.
-We evaluate robustness and detection performance under a diverse suite of white-box adversarial attacks, including Fast Gradient Sign Method (FGSM), Projected Gradient Descent (PGD), and Carlini–Wagner (CW) attacks.
-We also apply AutoAttack (standard version) and Square Attack [2], both of which are parameter-free (black-box) and ensemble-based.
-To assess robustness under this stronger threat model, we incorporate the state-of-the-art adaptive attack, Orthogonal-PGD [4], which optimizes adversarial perturbations not only to fool the classifier but also to minimize detection signals.
-This provides a rigorous evaluation of the proposed methods under any conditions.
-Implementation Details.
-All models are implemented in PyTorch and trained on NVIDIA RTX 6000 Ada.
-We use a batch size of 32, AdamW optimizer with learning rate $1\times 10^{-4}$ and weight decay of $0.01$, training the recover modules and augmentation matrices for 50 epochs.
-4.2 Detection Performance under Standard Attacks
-Table 2: The AUC of Different Adversarial Detection Approaches on CIFAR-10.
-The results are the mean and standard deviation of 5 runs.
-Our methods are included for comparison.
-Classifier: ResNet110, FGSM: $\epsilon=0.05$, PGD: $\epsilon=0.02$.
-Note that our methods and BEYOND need no AE for training, leading to the same value on both seen and unseen settings.
-The bolded values are the best performance, and the underlined italicized values are the second-best performance.
-Method Unseen (Attacks in training are excluded from tests) Seen (Attacks in training are included in tests)
-FGSM PGD AutoAttack Square FGSM PGD CW AutoAttack Square
-DkNN [27] 61.55 ±0.023 51.22 ±0.026 52.12 ±0.023 59.46 ±0.022 61.55 ±0.023 51.22 ±0.026 61.52 ±0.028 52.12 ±0.023 59.46 ±0.022
-kNN [9] 61.83 ±0.018 54.52 ±0.022 52.67 ±0.022 73.39 ±0.020 61.83 ±0.018 54.52 ±0.022 62.23 ±0.019 52.67 ±0.022 73.39 ±0.020
-LID [23] 71.08 ±0.024 61.33 ±0.025 55.56 ±0.021 66.18 ±0.025 73.61 ±0.020 67.98 ±0.020 55.68 ±0.021 56.33 ±0.024 85.94 ±0.018
-Hu [14] 84.51 ±0.025 58.59 ±0.028 53.55 ±0.029 95.82 ±0.020 84.51 ±0.025 58.59 ±0.028 91.02 ±0.022 53.55 ±0.029 95.82 ±0.020
-Mao [24] 95.33 ±0.012 82.61 ±0.016 81.95 ±0.020 85.76 ±0.019 95.33 ±0.012 82.61 ±0.016 83.10 ±0.018 81.95 ±0.020 85.76 ±0.019
-LNG [1] 98.51 63.14 58.47 94.71 99.88 91.39 89.74 84.03 98.82
-BEYOND [39] 98.89 ±0.013 99.28 ±0.020 99.16 ±0.021 99.27 ±0.016 98.89 ±0.013 99.28 ±0.020 99.20 ±0.008 99.16 ±0.021 99.27 ±0.016
-Our Approaches
-RT 99.93 ±0.005 96.89 ±0.071 $9$0.000 85.38 ±0.344 99.93 ±0.005 96.89 ±0.071 $9$0.002 $9$0.000 85.38 ±0.344
-LT 97.50 ±0.038 98.61 ±0.042 99.60 ±0.018 97.47 ±0.036 97.50 ±0.038 98.61 ±0.042 97.08 ±0.027 99.60 ±0.018 97.47 ±0.036
-RLT 99.85 ±0.005 99.37 ±0.011 $9$0.000 95.95 ±0.102 99.85 ±0.005 99.37 ±0.011 99.91 ±0.004 $9$0.000 95.95 ±0.102
-Table 3: The AUC of Different Adversarial Detection Approaches on ImageNet.
-To align with baselines, classifier: DenseNet121, FGSM: $\epsilon=0.05$, PGD: $\epsilon=0.02$.
-Due to memory and resource constraints, baseline methods are not evaluated against AutoAttack on ImageNet.
-Method Unseen Seen
-FGSM PGD FGSM PGD CW
-DkNN [27] 89.16 ±0.038 78.00 ±0.041 89.16 ±0.038 78.00 ±0.041 68.91 ±0.044
-kNN [9] 51.63 ±0.04 51.14 ±0.039 51.63 ±0.04 51.14 ±0.039 50.73 ±0.04
-LID [23] 90.32 ±0.046 52.56 ±0.038 99.24 ±0.043 98.09 ±0.042 58.83 ±0.041
-Hu [14] 72.56 ±0.037 86.00 ±0.042 72.56 ±0.037 86.00 ±0.042 80.79 ±0.044
-LNG [1] 96.85 89.61 99.53 98.42 86.05
-BEYOND [39] 97.59 ±0.04 96.26 ±0.045 97.59 ±0.04 96.26 ±0.045 95.46 ±0.047
-Our Approaches
-RT 94.31 ±0.457 $9$0.000 94.31 ±0.457 $9$0.000 92.18 ±0.135
-LT 96.18 ±0.028 97.89 ±0.021 96.18 ±0.028 97.89 ±0.021 94.06 ±0.215
-RLT 97.60 ±0.048 $9$0.000 97.60 ±0.048 $9$0.000 91.19 ±0.022
-We evaluate the effectiveness of our proposed detection scores, namely RT, LT, and RLT, on CIFAR-10 and ImageNet under standard adversarial threat models, including FGSM, PGD, CW, AutoAttack, and Square Attack.
-We benchmark against established baselines such as LID [23], DkNN [27], LNG [1], and the recent SSL-based BEYOND [39].
-The Area Under the Receiver Operating Characteristic Curve (RoC-AUC) results, which measure detection performance on various thresholds, are reported in Table 2 for CIFAR-10 and Table 3 for ImageNet.
-On CIFAR-10, RT achieves exceptionally strong performance across attacks that disrupt internal representations (e.g., FGSM, CW, AutoAttack), aligning with its role in capturing large, localized deviations in intermediate layers.
-LT performs particularly well against attacks like PGD and Square, which introduce minimal internal distortion but induce instability at the output layer.
-Notably, the fused RLT score achieves either the best or second-best AUC across most of the attacks, validating the complementary nature of RT and LT.
-Despite relying solely on the internal signals of the target classifier and requiring no adversarial examples or external models, RLT matches or outperforms state-of-the-art methods, including BEYOND, which uses large pre-trained SSL representations.
-On ImageNet, we observe similar trends.
-RT continues to perform well under strong perturbations such as PGD, while LT maintains robust accuracy under less structured noise like FGSM.
-The combined RLT score again leads to superior or competitive detection performance.
-Unlike several baselines that degrade substantially in the large-scale setting due to increased model capacity or overfitting to specific attacks, our method remains stable without requiring additional training resources or architecture-specific tuning.
-4.3 Detection Performance under Adaptive Attacks
-To assess the robustness of our proposed method against strong adaptive adversaries, we evaluate the RLT score under the Orthogonal-PGD adaptive attack, following the protocol in BEYOND [39], where the attacker is assumed to have full access to both the target classifier and the detection mechanism, and crafts perturbations specifically to evade detection by inducing misclassification.
-Table 4 reports the robust accuracy (RA) of various defense methods at two $L_{\infty}$ threat levels ($0.01$ and $8/255$), each evaluated at false positive rates (FPR) of 5% and 50%.
-While BEYOND achieves strong performance with the aid of additional components such as self-supervised models, our method achieves competitive performance using only the internal consistency of the target model, without requiring adversarial examples or external models during training.
-These results highlight the effectiveness of our RLT-based detection framework even in adversarially optimized threat settings.
-Table 4: RA(%) under Orthogonal-PGD Adaptive Attack using CIFAR-10 and ResNet110.
-Defense $\bm{L_{\infty}=0.01}$ $\bm{L_{\infty}=8/255}$
-RA@FPR5% RA@FPR50% RA@FPR5% RA@FPR50%
-Ours (RLT) 75.40 99.58 33.70 80.77
-BEYOND [39] 88.38 98.81 13.80 48.20
-Trapdoor [29] 0.00 7.00 0.00 8.00
-DLA [31] 62.60 83.70 0.00 28.20
-SID [32] 6.90 23.40 0.00 1.60
-SPAM [19] 1.20 46.00 0.00 38.00
-4.4 Empirical Evaluation of the Proposed Assumption
-Figure 3: Empirical validation of the A Few Large Shifts Assumption using CIFAR-10 and ResNet-110. We plot layer-wise error distributions $\sigma(e_{k})$ under different attack methods.
-To validate our A Few Large Shifts Assumption, we measured layer-wise reconstruction error distributions $\sigma(e_{k})$ on a ResNet-110 model trained on CIFAR-10 under various attacks: FGSM, PGD, CW, AutoAttack, and Square Attack.
-As shown in Figure 3, benign inputs exhibit relatively flat error profiles across layers, while certain attacks as shown in Figure 3.(a), produce sharp peaks at specific layers—indicating large indicated as red arrows, localized shifts in internal representations.
-These peaked patterns align with high RT detection performance and support the assumption that adversarial perturbations affect only a subset of layers $\mathcal{T}\subseteq\{f_{1},f_{2},\cdots,f_{L}\}$.
-In contrast, attacks like PGD and Square, which show stronger LT detection, yield smoother error curves similar to benign inputs but with greater variance as depicted in Figure 3.(b).
-This suggests that these attacks primarily affect the logit layer, with minimal disruption to intermediate features $\mathcal{T}=\{f_{logit}\}$.
-These findings empirically confirm that adversarial manipulations often concentrate at a few layers and further justify the complementary roles of RT and LT in our detection framework.
-4.5 Implementation Costs
-We assess the computational efficiency of our RLT method against existing baselines.
-As shown in Table 5, RLT introduces minimal floating-point operations (FLOPs), parameters, and model size overhead, offering a lightweight, plug-in solution without retraining or external models.
-It is significantly more efficient than SSL-based (Mao et al.[24]) and graph-based (LNG[1]) detectors, making it well-suited for real-time and resource-constrained deployments.
-Table 5: Comparison of implementation costs in terms of FLOPs, parameters, and model size overhead when the target model is ResNet110 with the CIFAR-10 dataset.
-Method FLOPs (G) Params (M) Model Size Overhead ($\bm{\times}$)
-Mao [24] 5.25 38.12 22.02
-LNG [1] 0.286 8.33 4.81
-BEYOND [39] 0.715 20.62 11.91
-Ours (RLT) 0.491 2.59 1.49
-5 Limitations and Future Work
-Our proposed framework uses RT and LT to detect adversarial examples by leveraging internal layer inconsistencies.
-Despite its simplicity, efficiency, and strong performance across diverse settings, it has several limitations that we now discuss.
-Our framework hinges on the A Few Large Shifts principle—that adversarial perturbations manifest predominantly in either a small subset of internal layers $\{f_{1},\dots,f_{L}\}$ or (separately) at the logit stage.
-While this binary decomposition keeps RT and LT analytically clean and easy to interpret, in practice highly adaptive or high-capacity attacks may simultaneously disrupt multiple layers and the logit outputs, creating overlapping distortions.
-In such mixed-mode scenarios, our current score fusion may underestimate these interactions, potentially reducing detection accuracy.
-Evaluating the detector under attacks designed to violate this assumption—and developing extensions to handle truly mixed shifts—constitutes important future work.
-Moreover, the logit space itself poses unique challenges for detection.
-Unlike intermediate feature maps, which admit spatial or structural comparisons, logits lack an interpretable geometric structure and offer no downstream transform against which to gauge stability.
-As a result, it is difficult to distinguish between benign variability—due to natural input perturbations or augmentations—and adversarial influence.
-Although our LT score partially mitigates this by using entropy-weighted changes under augmentations, it remains only a coarse proxy for true perturbation sensitivity.
-Finally, following prior work [39], we have evaluated our method only on CNNs (e.g., ResNet-110) and vision benchmarks.
-Extending our approach to modern architectures such as Transformers [33], as well as stress-testing it on assumption-violating and mixed-mode attacks, are important directions for future work.
-6 Conclusion
-In this work, we introduced a novel framework for adversarial example detection that leverages layer-wise inconsistencies within deep neural networks.
-Motivated by our proposed A Few Large Shifts Assumption–which posits that adversarial perturbations primarily cause substantial deviations in a small subset of network layers–we developed two complementary detection strategies: RT, capturing discrepancies in intermediate-layer reconstructions, and LT, quantifying output-layer sensitivity under minimal augmentations.
-Extensive evaluations on CIFAR-10, CIFAR-100, and ImageNet demonstrate that our combined approach, RLT achieves state-of-the-art detection performance, significantly outperforming existing baselines while remaining highly robust to adaptive adversarial attacks.
-Our method operates efficiently without relying on external models, adversarial training, or extensive augmentation, highlighting its practical suitability for real-world deployment.
-Acknowledgements
-This work was supported in part by the DARPA Young Faculty Award, the National Science Foundation (NSF) under Grants #2127780, #2319198, #2321840, #2312517, and #2235472, the Semiconductor Research Corporation (SRC), the Office of Naval Research through the Young Investigator Program Award, and Grants #N00014-21-1-2225 and #N00014-22-1-2067, Army Research Office Grant #W911NF2410360, and the U.S.
-Army Combat Capabilities Development Command (DEVCOM) Army Research Laboratory under Support Agreement No.
-USMA 21050.
-Additionally, support was provided by the Air Force Office of Scientific Research under Award #FA9550-22-1-0253, along with generous gifts from Xilinx and Cisco.
-The views expressed in this paper are those of the authors and do not reflect the official policy or position of the U.S.
-Military Academy, the U.S.
-Army, the U.S.
-Department of Defense, or the U.S.
-Government.
-Appendix A Theoretical Analysis
+In contrast to the methods above, our framework introduces a fundamentally different detection paradigm.
+Conceptualizing the network as a cylinder where each layer is a horizontal slice, prior work can be seen as performing a “horizontal” analysis: comparing a sample’s features against a reference set or augmentations within a single layer.
+We propose a “vertical” analysis that is entirely self-referential.
+Our method scrutinizes the consistency of feature transformations between the network’s own layers for a single input, eliminating the need for the external reference data, heavy augmentations, or complex data structures common in horizontal methods.
+This approach is grounded in our A Few Large Shifts Assumption, which posits that adversarial perturbations cause large, localized deviations between a few critical layers, making these vertical inconsistencies a detectable signal.
+Appendix B Theoretical Analysis
 We provide formal justification for our proposed detection framework, which includes RT, LT, and the fused score RLT.
@@
 This justifies the detection reliability of our method under common threat models.
-A.1 Justification for Layer-wise Reconstruction Error in RT
+B.1 Connection to Lipschitz Continuity
+Our detection framework can be formally understood as an empirical method for identifying violations of Lipschitz continuity across localized layers, we term the violation of layer-wise Lipschitz continuity.
+Let $f_{k}$ be the function corresponding to the $k$-th layer of the network, such that $\mathbf{z}_{k}=f_{k}(\mathbf{z}_{k-1})$.
+The local Lipschitz constant $L_{k}$ at an input $\mathbf{z}_{k-1}$ is the smallest value such that:
+$\|f_{k}(\mathbf{z}^{\prime}_{k-1})-f_{k}(\mathbf{z}_{k-1})\|\leq L_{k}\|\mathbf{z}^{\prime}_{k-1}-\mathbf{z}_{k-1}\|$ (7)
+for all $\mathbf{z}^{\prime}_{k-1}$ in a neighborhood of $\mathbf{z}_{k-1}$.
+Our A Few Large Shifts Assumption posits that for an adversarial input $x^{\mathrm{adv}}$, the ratio $\frac{\|\mathbf{z}_{k+1}(x^{\mathrm{adv}})-\mathbf{z}_{k+1}(x)\|_{2}}{\|\mathbf{z}_{k}(x^{\mathrm{adv}})-\mathbf{z}_{k}(x)\|_{2}}$ becomes very large for a small subset of layers $\mathcal{T}$.
+This ratio serves as an empirical estimate of the local Lipschitz constant for the layer transformation $f_{k+1}$.
+Our Recovery Testing (RT) score is designed to detect these violations.
+A large reconstruction error, $e_{k}(x)$, signals a significant deviation between the expected and actual feature transformations, which is a manifestation of this high layer-wise Lipschitz constant.
+By calibrating the expected error distributions on benign data—where we assume layer-wise Lipschitz constants are small and stable—RT effectively identifies inputs that cause these localized instabilities.
+B.2 Justification for Layer-wise Reconstruction Error in RT
 This section analyzes why using the reconstruction error of intermediate features—as done in RT—produces a separable distribution between benign and adversarial inputs.
@@
 $\|z_{i}(x^{\mathrm{adv}})-R^{(L\rightarrow i)}(z_{L}(x^{\mathrm{adv}}))\|^{2}>\|z_{i}(x)-R^{(L\rightarrow i)}(z_{L}(x))\|^{2}.$
-(Proof: Appendix B.1)
-A.2 Justification for Ratio-based Logit-layer Deviation in LT
+(Proof: Appendix C.1)
+B.3 Justification for Ratio-based Logit-layer Deviation in LT
 In this section, we justify the use of the LT score, which captures how the logit-layer output changes under augmentation relative to changes in intermediate-layer features.
@@
 $\|\textbf{o}(\ell(x^{\mathrm{adv}}))-\textbf{o}(\ell(x^{(g),\text{adv}}))\|>0.$
-(Proof: Appendix B.2)
+(Proof: Appendix C.2)
 Theorem 3 (Amplified Logit Sensitivity).
@@
 $\frac{\|\ell^{(g),\text{adv}}-\ell^{\text{adv}}\|}{\|z^{(g),\text{adv}}-z^{\text{adv}}\|}>\frac{\|\ell^{(g)}-\ell\|}{\|z^{(g)}-z\|}.$
-(Proof: Appendix B.3)
-A.3 Quantile-normalized RT + LT Provides Jointly Separable Score in RLT
+(Proof: Appendix C.3)
+B.4 Quantile-normalized RT + LT Provides Jointly Separable Score in RLT
 We now provide a theoretical justification for the fused detection score $RLT(x)=RT_{\text{norm}}^{2}+LT_{\text{norm}}^{2}$.
@@
 Assumption 7 (Quantile-normalized RT and LT).
-Let ^$\hat{\mathcal{F}}_{RT},\hat{\mathcal{F}}_{LT}$ be empirical CDFs computed on benign RT and LT scores. Define:
+Let $\hat{\mathcal{F}}_{RT},\hat{\mathcal{F}}_{LT}$ be empirical CDFs computed on benign RT and LT scores. Define:
 $RT_{\text{norm}}(x)=\Phi^{-1}(\hat{\mathcal{F}}_{RT}(RT(x))),\quad LT_{\text{norm}}(x)=\Phi^{-1}(\hat{\mathcal{F}}_{LT}(LT(x))),$
@@
 $\mathbb{E}[RLT(x)]=2\quad\text{for benign }x,\quad\text{and}\quad RLT(x^{\text{adv}})>\gamma^{2}.$
-(Proof : Appendix B.4)
-A.4 Robustness to Adaptive Attacks
+(Proof : Appendix C.4)
+B.5 Robustness to Adaptive Attacks
 In this section, we provide a theoretical perspective on the robustness of our detection framework under adaptive adversaries.
@@
 We show that the design of our RT and LT metrics inherently introduces conflicting optimization gradients, which hinder the adversary’s ability to jointly suppress both detection scores.
-A.4.1 Adaptive Attack Objective
+B.5.1 Adaptive Attack Objective
 To simulate a strong white-box adversary, we formulate an adaptive objective that simultaneously minimizes the classification loss and both detection scores (RT and LT).
@@
 The optimization becomes increasingly difficult when the gradients corresponding to RT and LT conflict in direction.
-A.4.2 Conflicting Gradient Effects
+B.5.2 Conflicting Gradient Effects
 The RT score is defined as:
@@
 In practice, we observe that attempting to suppress one score often exacerbates the other, making it difficult to minimize the fused detection score $RLT(x)=RT_{\text{norm}}^{2}(x)+LT_{\text{norm}}^{2}(x)$.
-Appendix B Proofs
-B.1 Theorem 1
+Appendix C Proofs
+C.1 Theorem 1
 Proof.
@@
 Squaring both sides proves the theorem. ∎
-B.2 Theorem 2
+C.2 Theorem 2
 Proof.
@@
 ∎
-B.3 Theorem 3
+C.3 Theorem 3
 Proof.
@@
 Hence, $\ell^{(g),\text{adv}}\approx\ell^{\text{adv}}$, contradicting Theorem 2. Thus the adversarial logit must be more sensitive to benign augmentation than the original. ∎
-B.4 Theorem 4
+C.4 Theorem 4
 Proof.
@@
 $RLT(x^{\text{adv}})>\gamma^{2}.$
-Appendix C Ablation Study
+Appendix D Ablation Study
 All ablation studies were conducted using the CIFAR-10 dataset with ResNet110. We systematically vary the key hyperparameters and architectural choices to assess their impact on detection performance.
-C.1 Size of recover module
+D.1 Size of recover module
 We evaluate the impact of varying the depth and dimensionality of the recover modules on RT detection performance.
@@
 512 0.00 0.00 0.00 0.00 0.00 0.00
-C.2 Number of learnable augmentations $G$
+D.2 Number of learnable augmentations $G$
 We investigate how varying the number of learnable augmentation matrices ($G$) impacts LT detection performance.
@@
 4 0.00 0.00 0.00 0.00 0.00 0.00
-C.3 Choice of $k_{RT}$ and $k_{LT}$
+D.3 Choice of $k_{RT}$ and $k_{LT}$
 We examine how choosing identical values for hyperparameters $k_{RT}$ and $k_{LT}$ affects combined RLT performance relative to optimal, independent selection of these parameters.
@@
 $k_{RT}\neq k_{LT}$ (Optimal) 0.00 0.00 0.00 0.00 0.00 0.00
-C.4 Single objective adaptive attacks
+D.4 Single objective adaptive attacks
+To highlight the inherent robustness of our framework, we begin with an ablation study under a simplified setting.
+Specifically, we consider cases where both the attacker and the defender rely on a single detection objective—either RT or LT.
+As shown in Table 12, this restriction leads to substantial performance degradation compared to the full defense, where both RT and LT are jointly employed for attack and detection.
+These results demonstrate that neither component alone is sufficient.
 Table 12: Single Objective Orthogonal-PGD Adaptive Attack on $L_{\infty}=8/255$ using CIFAR-10 and ResNet110.
@@
 LT 17.20% 56.17%
-To further elucidate the inherent robustness of our detection framework, we conduct an ablation analysis by removing individual objectives from the Orthogonal-PGD adaptive attack.
-Specifically, we investigate scenarios where only one detection objective–either RT or LT–is targeted by the adversary.
-The results, summarized in Table 9, highlight significant performance drops when either RT or LT objectives are excluded.
-This empirical outcome clearly substantiates our theoretical analysis presented earlier: adaptive attacks face inherent difficulties due to conflicting gradient directions induced by RT and LT, which makes attack algorithms difficult to find $\mathcal{T}$.
-The interpretation of this phenomenon aligns precisely with our conceptual framework.
-When RT is removed, adversarial perturbations predominantly shift their targeted layers $\mathcal{T}$ towards intermediate layers, significantly affecting internal representations.
-Conversely, when LT is excluded, adversarial perturbations primarily influence the logit layer, creating instability at the output level.
-These distinct perturbation strategies confirm our assumption that RT and LT effectively impose opposing constraints on adaptive adversaries, complicating their optimization efforts to simultaneously minimize both scores.
-Remarkably, even under this weakened adaptive setting–where adversaries optimize only against one detection objective–our combined detection (RLT) continues to outperform current state-of-the-art baseline models, as reported in Table 4.
-This robust performance demonstrates that our quantification of layer-by-layer perturbations effectively captures adversarial manipulation, affirming the design efficacy of our proposed RT and LT measures in practical adversarial settings.
-C.5 Ablation on perturbation budget sensitivity of each testing measure
+To provide a more comprehensive analysis, we conducted an additional experiment where the attacker’s objective may differ from the defender’s measurement score.
+As shown in Table 13, an adversary attacking the full RLT objective achieves the lowest average Robust Accuracy (RA) across all defender configurations (64.85%).
+This demonstrates that the fused RLT framework poses a significant challenge for an adversary.
+The most effective evasion strategy is to target the combined RLT score directly, as attacking the individual RT or LT components results in a lower expected attack success rate.
+Table 13: RA@FPR50% (%) under mismatched end-to-end PGD attack and defense objectives ($\epsilon=8/255$, ResNet110, CIFAR-10).
+Measured With $\downarrow$ / Attacked Measure $\rightarrow$ RLT RT LT
+RLT $\mathbf{68.53_{\pm 2.21}}$ $72.06_{\pm 4.05}$ $73.05_{\pm 2.99}$
+RT $78.75_{\pm 1.53}$ $\mathbf{75.56_{\pm 1.30}}$ $81.58_{\pm 1.73}$
+LT $47.28_{\pm 4.43}$ $64.49_{\pm 8.12}$ $\mathbf{46.48_{\pm 3.67}}$
+Average 64.85 70.70 67.04
+This outcome empirically substantiates our theoretical analysis: adaptive attacks face inherent difficulties due to conflicting gradient directions induced by RT and LT, which makes attack algorithms difficult to find $\mathcal{T}$.
+When an adversary focuses only on evading RT, they tend to create perturbations that shift the logit layer, which are then caught by LT.
+Conversely, attacks targeting only LT tend to disrupt intermediate features, which are then caught by RT.
+These distinct perturbation strategies confirm our assumption that RT and LT effectively impose opposing constraints on adaptive adversaries, making the combined RLT score a significantly more robust defense.
+D.5 Ablation on perturbation budget sensitivity of each testing measure
 Figure 4: Ablation study showing AUC (%) of RT, LT, and RLT under varying perturbation budgets ($\epsilon\in\{0.01,0.02,\cdots,0.1\}$) across multiple standard attack types (FGSM, PGD, AutoAttack, Square).
@@
 The results further demonstrate that combining RT and LT provides stable and effective detection across a wide perturbation spectrum.
-C.6 Contribution of each term in testing measures
+D.6 Contribution of each term in testing measures
 Figure 5: Empirical distributions of individual terms used in our detection metrics, evaluated on CIFAR-10 with a ResNet-110 model under FGSM attack ($\epsilon=0.05$).
@@
 The final integration of both methods in RLT leads to the strongest separation, illustrating the complementary nature of these features.
-Appendix D AEs Detection Across Architectures on CIFAR-100
-Table 11: AUC scores (%) on CIFAR-100 under three white-box attacks (FGSM, $\ell_{\infty}$-PGD, and AutoAttack) with the same perturbation budget as the previous evaluations.
-ResNet-18 MobileNet-V2 x0 5
+D.7 Evaluation Against $L_{2}$-Norm Attacks
+To ensure our method’s effectiveness is not limited to a single threat model, we evaluated its performance against $L_{2}$-norm attacks, demonstrating that its detection capabilities are largely norm-agnostic.
+First, to establish a challenging and comparable evaluation setting, we identified a worst-case perturbation budget for an $L_{2}$-PGD attack.
+As shown in Table 15, a budget of $\epsilon=0.6$ was sufficient to reduce the base classifier’s Robust Accuracy to 0.00%.
+Table 15: RA (%) under $L_{2}$-PGD attack on CIFAR-10 to determine a worst-case budget.
+$\bm{\epsilon}$ RA
+0.1 39.97
+0.2 6.21
+0.3 0.74
+0.4 0.10
+0.5 0.01
+0.6 0.00
+Using this budget, we then compared our detector’s performance against both $L_{\infty}$ and $L_{2}$ attacks.
+The results in Table 16 show that our combined RLT detector achieved a high AUC of 98.05% against the strong $L_{2}$ attack, which is comparable to its performance against the $L_{\infty}$ attack.
+This confirms the robustness of our detection framework across different norm-based threat models.
+Table 16: AUC (%) scores comparing performance against $L_{\infty}$ and $L_{2}$ PGD attacks on CIFAR-10.
+Norm RA RLT RT LT
+$L_{\infty}$ 0.00 99.47 99.49 96.65
+$L_{2}$ 0.00 98.05 95.05 97.16
+Appendix E Generalization to CNN-based Architectures
+Table 17: AUC scores (%) on CIFAR-100 under three white‑box attacks (FGSM, $\ell_{\infty}$‑PGD, and AutoAttack) with the same perturbation budget as the previous evaluations.
+ResNet‑18 MobileNet‑V2 x0 5
 Variant FGSM PGD AutoAttack Avg. FGSM PGD AutoAttack Avg.
@@
 RLT 95.61 98.92 99.97 98.17 98.73 98.14 99.98 98.95
-ShuffleNet-V2 x0 5 RepVGG a0
+ShuffleNet‑V2 x0 5 RepVGG a0
 RT 98.38 97.35 99.97 98.57 91.77 98.94 99.99 96.90
@@
 RLT 98.16 96.17 99.94 98.09 93.49 98.46 99.98 97.31
-To assess the generalizability of our detection framework across diverse neural architectures, we evaluate RT, LT, and RLT on CIFAR-100 using four distinct backbone models: ResNet-18, MobileNet-V2 x0 5, ShuffleNet-V2 x0 5, and RepVGG-a0.
+To assess the generalizability of our detection framework across diverse neural architectures, we evaluate RT, LT, and RLT on CIFAR-100 using four distinct backbone models: ResNet‑18, MobileNet‑V2 x0 5, ShuffleNet‑V2 x0 5, and RepVGG-a0.
 As shown in Table 17, our methods consistently achieve high AUC scores across all architectures and attack types (FGSM, PGD, and AutoAttack), with RLT offering the most robust overall performance.
@@
 This plug-and-play nature, combined with high detection efficacy, makes our method highly suitable for practical applications where architectural diversity and efficiency are key considerations.
+Appendix F Generalization to Transformer-based Architecture
+To evaluate the generality of our detection framework beyond CNNs, we applied it to a pre-trained Vision Transformer (ViT-B/16) on CIFAR-10 and CIFAR-100. The results are presented in Table 18.
+[…]
+As shown in the table, our method consistently outperforms the LID baseline on CIFAR-10.
+On CIFAR-100, it matches or exceeds LID’s performance against PGD and CW attacks while remaining competitive against FGSM.
+These strong results, achieved with minimal overhead, confirm that our layer-wise inconsistency-based detection strategy generalizes effectively to modern Transformer architectures and offers significant potential for further performance tuning.
+Table 18: AUC (%) scores for ViT-B/16 on CIFAR-10 and CIFAR-100 against various attacks ($\epsilon=0.03$).
+Dataset Method FGSM PGD CW
+CIFAR-10 LID 92.65 82.89 67.90
+Ours (RLT) 95.14 90.68 99.99
+CIFAR-100 LID 91.05 81.28 74.37
+Ours (RLT) 88.18 84.54 99.99
+Appendix G Detailed Computational Cost Analysis
+To substantiate our claim of low computational overhead, we provide a detailed analysis comparing our framework to two common detection paradigms: SSL-based and reference-set-based detectors.
+The analysis uses the target models detailed in Table 19.
+[…]
+Second, reference-set detectors incur significant memory costs for storing embeddings and, in the case of graph-based methods, adjacency matrices.
+As detailed in Table 21, this overhead scales with the size of the reference set and can be prohibitively large, especially for graph-based approaches.
+[…]
+Table 19: Target models used in our experiments, along with their parameter counts and the number of feature blocks (layers) used for applying RT.
+Target Model # Parameters # Feature Blocks
+MobileNet 815,780 19
+ShuffleNet 1,356,104 18
+ResNet110 1,730,714 56
+RepVGG 7,956,164 23
+Table 20: Evaluation of SSL-based approaches’ minimum overhead ($\times$) introduced by pre-trained SSL models relative to each target model.
+SSL Models # Params MobileNet ShuffleNet ResNet110 RepVGG
+BYOL 25,557,032 31.33 18.85 14.77 3.21
+SimSiam 38,201,408 46.83 28.17 22.07 4.80
+MoCo v3 (ViT) 215,678,464 264.38 159.04 124.62 27.11
+Table 21: Minimum overhead ($\times$) from detection approaches using a reference set with a 1024-dimensional embedding space.
+"Graph Structure" indicates whether the approach constructs an adjacency matrix.
+Graph Structure Reference Set Size # Params MobileNet ShuffleNet ResNet110 RepVGG
+No 1,000 1,024,000+ 1.26+ 0.76+ 0.59+ 0.13+
+5,000 5,120,000+ 6.28+ 3.78+ 2.96+ 0.64+
+40,000 40,960,000+ 50.21+ 30.20+ 23.67+ 5.15+
+Yes 1,000 2,024,000+ 2.48+ 1.49+ 1.17+ 0.25+
+5,000 30,120,000+ 36.92+ 22.21+ 17.40+ 3.79+
+40,000 1,640,960,000+ 2011.52+ 1210.05+ 948.14+ 206.25+
+Table 22: Overhead ($\times$) and AUC performance (%) (CIFAR-100 under $\ell_{\infty}$‑PGD attack with $\epsilon=0.02$) of our approach across different target models and varying depth and width of the recovery modules.
+"Performance Loss" indicates the AUC drop relative to the largest detector configuration for the same target model.
+Target Model Depth Width Detector’s # Params Overhead AUC Performance Loss
+MobileNet 2 64 1,540,744 1.89 97.37 0.0000
+2 32 776,904 0.95 97.20 -0.1665
+2 16 394,984 0.48 96.07 -1.3025
+2 8 204,024 0.25 95.12 -2.2492
+2 4 108,544 0.13 89.27 -8.0933
+2 2 60,804 0.07 87.66 -9.7059
+ShuffleNet 2 128 2,788,840 2.06 96.07 0.0000
+4 64 1,548,456 1.14 95.46 -0.6154
+2 64 1,402,664 1.03 95.96 -0.1152
+4 32 747,656 0.55 95.12 -0.9522
+2 32 709,576 0.52 95.48 -0.5946
+4 16 373,368 0.28 93.88 -2.1923
+2 16 363,032 0.27 95.25 -0.8218
+4 8 192,752 0.14 93.02 -3.0472
+2 8 189,760 0.14 94.56 -1.5079
+ResNet110 3 256 5,160,992 2.98 99.84 0.0000
+2 256 1,514,272 0.87 99.58 -0.2640
+3 64 652,448 0.38 99.71 -0.1259
+2 64 416,608 0.24 99.57 -0.2717
+3 16 158,912 0.09 99.51 -0.3313
+2 16 142,192 0.08 99.41 -0.4252
+3 4 75,128 0.04 99.62 -0.2180
+2 4 73,588 0.04 99.63 -0.2119
+RepVGG 4 256 11,310,992 1.42 98.45 0.0000
+3 256 9,852,304 1.24 98.47 0.0144
+4 128 4,942,992 0.62 98.38 -0.0721
+3 128 4,574,096 0.57 98.41 -0.0386
+4 64 2,299,664 0.29 98.34 -0.1092
+3 64 2,205,328 0.28 98.36 -0.0969
+4 32 1,113,168 0.14 98.23 -0.2185
+3 32 1,088,528 0.14 98.32 -0.1341
+4 16 553,712 0.07 98.01 -0.4442
+3 16 547,024 0.07 98.11 -0.3468
+Appendix H Detailed Plug-in-play System-Level Analysis
+In this section, we provide a detailed theoretical and empirical analysis of the system-level performance when our detector is applied as a plug-in module to a standard classifier.
+We formally define the evaluation metrics, introduce a novel framework for establishing guaranteed lower bounds on system accuracy, and demonstrate how this framework can be used to select optimal detection thresholds in practical scenarios.
+H.1 Definitions of Metrics
+To formally analyze performance, we define metrics for both the base classifier and the combined classifier-detector system.
+Let $C(x)$ be the classifier’s prediction for an input $x$, $y$ be its true label, and $D(x)$ be our detector’s output, where $D(x)=1$ signifies an adversarial detection.
+  • Classifier Clean Accuracy ($CA_{cls}$): The accuracy of the base classifier on benign samples without any detector. $CA_{cls}=\mathbb{P}(C(x_{\text{clean}})=y)$.
+  • Classifier Robust Accuracy ($RA_{cls}$): The accuracy of the base classifier on adversarial samples without any detector. $RA_{cls}=\mathbb{P}(C(x_{\text{adv}})=y)$.
+  • System Clean Accuracy ($CA_{sys}$): The accuracy of the combined system on benign samples.
+A benign sample is handled correctly only if it is both correctly classified and not flagged by the detector.
+    $CA_{sys}=\mathbb{E}[\mathbb{I}(C(x_{\text{clean}})=y\land D(x_{\text{clean}})=0)]$
+  • System Robust Accuracy ($RA_{sys}$): The accuracy of the combined system on adversarial samples.
+An adversarial sample is successfully defended if it is either detected or correctly classified despite the attack.
+    $RA_{sys}=\mathbb{E}[\mathbb{I}(C(x_{\text{adv}})=y\lor D(x_{\text{adv}})=1)]$
+  • Overall System Accuracy ($A_{sys}$): The expected accuracy of the system given a probability $p$ that an input is adversarial.
+    $A_{sys}=(1-p)\cdot CA_{sys}+p\cdot RA_{sys}$
+  • False Positive Rate (FPR): The fraction of benign samples incorrectly flagged as adversarial. $FPR=\mathbb{P}(D(x_{\text{clean}})=1)$.
+  • True Positive Rate (TPR): The fraction of adversarial samples correctly flagged as adversarial. $TPR=\mathbb{P}(D(x_{\text{adv}})=1)$.
+H.2 Performance Analysis Across Operating Points
+While AUC provides an aggregate measure, evaluating performance at fixed operating points is critical for understanding the practical trade-off between clean accuracy and robustness.
+First, we establish the baseline performance of the target ResNet-110 classifier without our defense in Table 23.
+The results show that while the model achieves high clean accuracy, its robustness is completely compromised by strong attacks like PGD and AutoAttack, with the $RA_{cls}$ dropping to 0.00%.
+Table 23: Baseline CA and RA (%) of the undefended ResNet-110 on CIFAR-10 ($\epsilon=8/255$).
+Attack $CA_{cls}$ $RA_{cls}$
+FGSM 92.49 25.75
+PGD 92.49 0.00
+CW 92.49 47.17
+AutoAttack 92.49 0.00
+In contrast, with our RLT detector active, the system’s performance is drastically improved, as detailed in Table 24.
+At a modest 5% FPR, the system maintains a high $CA_{sys}$ of 88.70% while restoring the $RA_{sys}$ against PGD from 0% to 99.27%.
+The table further illustrates the clear trade-off available to a practitioner: increasing the FPR boosts the TPR and, consequently, the $RA_{sys}$, at the cost of $CA_{sys}$.
+Table 24: Measured TPR, System Clean Accuracy ($CA_{sys}$), and System Robust Accuracy ($RA_{sys}$) (%) at varying FPRs for the defended system on CIFAR-10 ($\epsilon=8/255$).
+Attack Metric @FPR5% @FPR10% @FPR25% @FPR50%
+FGSM TPR 84.84 94.13 99.21 99.94
+$CA_{sys}$ 88.71 84.52 71.60 48.80
+$RA_{sys}$ 90.80 96.57 99.58 99.96
+PGD TPR 99.27 99.68 99.92 99.95
+$CA_{sys}$ 88.70 84.73 71.71 48.70
+$RA_{sys}$ 99.27 99.68 99.92 99.95
+CW TPR 90.46 96.48 99.48 99.97
+$CA_{sys}$ 88.63 84.57 71.79 48.77
+$RA_{sys}$ 96.14 98.79 99.91 100.00
+AutoAttack TPR 88.86 89.90 91.88 93.87
+$CA_{sys}$ 88.70 84.73 71.71 48.70
+$RA_{sys}$ 88.86 89.90 91.88 93.87
+H.3 Plug-in Robustness Gains with Adversarially Trained Classifiers (ATC)
+A key advantage of our detection framework is its role as a modular, plug-in defense.
+Rather than replacing robust training methods, our detector can be integrated with existing robust models, such as Adversarially Trained Classifiers (ATCs), to further enhance their performance.
+The value of a detector in this context is measured by the robustness improvement it provides to the overall system.
+H.3.1 Theoretical Justification for Robustness Improvement
+The performance of a combined classifier-detector system can be formally analyzed.
+The expected end-to-end system robust accuracy, $RA_{sys}$, is a function of the detector’s TPR and the base classifier’s own robust accuracy ($RA_{cls}$).
+The expected system robustness is given by:
+$RA_{sys}\approx TPR+(1-TPR)\times RA_{cls}$ (8)
+This can be rewritten as:
+$RA_{sys}\approx RA_{cls}+TPR\times(1-RA_{cls})$ (9)
+This relationship mathematically demonstrates that any detector with a non-zero TPR ($TPR>0$) is expected to improve the system’s robustness over the base classifier alone ($RA_{sys}>RA_{cls}$).
+H.3.2 Empirical Validation
+To empirically validate this theoretical relationship, we applied our detector to several ATCs with varying levels of baseline robustness. The results, presented in Table 25, confirm two key points:
+  • The measured $RA_{sys}$ shows a substantial improvement over the initial $RA_{cls}$ in all cases.
+For instance, a classifier with a baseline robustness of 55.88% achieves a final robustness of 83.08% when paired with our detector.
+  • The measured $RA_{sys}$ values closely correspond to the expected values predicted by our formula, validating its utility as a model for system performance and confirming the plug-in value of our detector.
+Table 25: Expected vs.
+Measured System Robust Accuracy ($RA_{sys}$) (%) when applying our detector to Adversarially Trained Classifiers with varying baseline robustness ($RA_{cls}$) under an adaptive attack.
+$RA_{cls}$ (%) Detector TPR (%) Expected $RA_{sys}$ (%) Measured $RA_{sys}$ (%) Improvement $\Delta$ (%)
+6.84 72.13 74.03 77.44 +70.6
+12.71 63.74 68.34 74.57 +61.86
+55.88 73.03 88.10 83.08 +27.2
+H.4 Formal System Accuracy Guarantees with Lower Bounds
+General Lower Bound. When the attack probability $p$ is unknown, we can establish a general lower bound on system accuracy.
+Theorem 5.
+Let $p,CA_{sys},RA_{sys}\in[0,1]$. Then $(1-p)CA_{sys}+p\cdot RA_{sys}\geq CA_{sys}\cdot RA_{sys}$.
+Proof.
+Let $f(p)=(1-p)CA_{sys}+p\cdot RA_{sys}-CA_{sys}\cdot RA_{sys}$.
+As a linear function of $p$ over the interval $[0,1]$, its minimum must occur at an endpoint.
+At $p=0$, $f(0)=CA_{sys}(1-RA_{sys})\geq 0$.
+At $p=1$, $f(1)=RA_{sys}(1-CA_{sys})\geq 0$.
+Since the function is non-negative at both endpoints, the inequality holds for all $p\in[0,1]$.
+∎
+Adaptive Lower Bound. For practical scenarios where we can assume an upper bound on the attack probability ($p\leq p^{\prime}$), we can derive a tighter lower bound.
+Theorem 6.
+Let $p\in[0,p^{\prime}]$ and $CA_{sys},RA_{sys},p^{\prime}\in[0,1]$. Then the system accuracy is lower-bounded by:
+$(1-p)CA_{sys}+p\cdot RA_{sys}\geq CA_{sys}\cdot RA_{sys}+\max(0,(1-p^{\prime})(CA_{sys}-RA_{sys}))$
+Proof.
+From Theorem 1, for any $t\in[0,p^{\prime}]$, we have $(1-t)CA_{sys}+t\cdot RA_{sys}\geq CA_{sys}\cdot RA_{sys}$.
+Let $t=p^{\prime}-p$.
+Since $p\in[0,p^{\prime}]$, $t$ is also in $0,p^{\prime}]$.
+Substituting $t=p^{\prime}-p$ yields:
+$(1-p)CA_{sys}+p\cdot RA_{sys}\geq CA_{sys}\cdot RA_{sys}+(1-p^{\prime})(CA_{sys}-RA_{sys})$
+This bound is tighter than the general one only when $CA_{sys}>RA_{sys}$. We therefore take the maximum of the additional term and zero to ensure the tightest possible bound in all cases.
+$(1-p)CA_{sys}+p\cdot RA_{sys}\geq CA_{sys}\cdot RA_{sys}+\max(0,(1-p^{\prime})(CA_{sys}-RA_{sys}))$
+H.5 Empirical Validation and Optimal Threshold Selection
+This theoretical framework provides a rigorous method for selecting an optimal operating point.
+We applied this analysis to our FGSM attack results in Table 26.
+Maximizing the general lower bound ($CA_{sys}\times RA_{sys}$) suggests an FPR of 10% is optimal.
+However, by maximizing the Adaptive Lower Bound for a realistic low attack probability (e.g., $p^{\prime}\leq 1\%$), we find that a 1% FPR provides a better performance guarantee.
+This confirms that our framework allows for principled configuration of the defense based on practical deployment assumptions.
+Table 26: Analysis of system performance and lower bounds (%) at different FPRs for the FGSM attack ($\epsilon=8/255$).
+FPR TPR $CA_{sys}$ $RA_{sys}$ Lower Bound Est. Lower Bound $\bm{A_{sys}}$ for attack prob. $\bm{p}$ Adaptive Lower Bound for max prob. $\bm{p^{\prime}}$
+0.1% 1% 5% 10% 50% 90% 95% 99% 0.1% 1% 5% 10% 50% 90% 95% 99%
+1 57.67 91.77 71.43 67.05 65.55 91.75 91.57 90.75 89.74 81.60 73.46 72.45 71.63 85.87 85.69 84.87 83.86 75.72 67.59 66.57 65.75
+5 86.79 88.53 91.54 81.59 81.04 88.53 88.56 88.68 88.83 90.04 91.24 91.39 91.51 81.04 81.04 81.04 81.04 81.04 81.04 81.04 81.04
+10 95.42 84.60 97.26 82.53 82.28 84.61 84.73 85.23 85.87 90.93 95.99 96.63 97.13 82.28 82.28 82.28 82.28 82.28 82.28 82.28 82.28
+15 97.75 80.34 98.64 79.41 79.25 80.36 80.52 81.25 82.17 89.49 96.81 97.73 98.46 79.25 79.25 79.25 79.25 79.25 79.25 79.25 79.25
+20 99.00 76.07 99.42 75.64 75.63 76.09 76.30 77.24 78.40 87.74 97.08 98.25 99.19 75.63 75.63 75.63 75.63 75.63 75.63 75.63 75.63
+25 99.34 71.66 99.63 71.37 71.39 71.69 71.94 73.06 74.46 85.64 96.83 98.23 99.35 71.39 71.39 71.39 71.39 71.39 71.39 71.39 71.39
+30 99.63 67.21 99.80 67.08 67.08 67.24 67.54 68.84 70.47 83.51 96.54 98.17 99.47 67.08 67.08 67.08 67.08 67.08 67.08 67.08 67.08
+Appendix I Robustness to Benign Noise
+To assess our detector’s specificity and ensure it is not merely flagging any large perturbation, we evaluated its response to significant yet benign noise.
+We conducted an experiment by applying random noise of varying magnitudes ($\epsilon$) to the unseen test set.
+Crucially, to isolate the effect of the perturbation itself from a label change, we only kept noise instances that were “benign” in their outcome, meaning they did not alter the classifier’s original prediction.
+We then measured the new False Positive Rate (FPR) using thresholds that were originally calibrated on clean, unperturbed data.
+The findings, presented in Table 27, provide strong evidence that our detector is not simply flagging any large perturbation but is specifically sensitive to the structure of adversarial attacks.
+At low-to-moderate noise levels ($\epsilon=4/255$ and $8/255$), the detector remained highly stable, with the FPR remaining nearly unchanged.
+Even with substantial random noise ($\epsilon=32/255$), the detector’s response was moderate; for example, a threshold calibrated for 5% FPR on clean data resulted in a new FPR of only 17.51%.
+This suggests that the large latent shifts our method identifies are a characteristic feature of crafted, adversarial perturbations.
+Table 27: Measured False Positive Rate (FPR) (%) on Benign Data with Random Noise.
+This table shows the new FPR when applying thresholds that were originally calibrated to give 1%, 5%, 10%, etc., FPR on clean data.
+The test is repeated for different magnitudes ($\epsilon$) of benign random noise.
+$\bm{\epsilon}$ FPR@1% FPR@5% FPR@10% FPR@15% FPR@20% FPR@25% FPR@30% FPR@35% FPR@40% FPR@45% FPR@50%
+4/255 0.93 4.77 9.27 13.60 18.51 23.03 27.61 32.01 36.27 41.20 46.26
+8/255 1.05 4.46 8.04 11.79 15.65 19.60 23.52 27.65 32.07 36.61 41.15
+32/255 12.38 17.51 22.16 26.71 31.39 35.99 40.34 44.72 49.16 53.54 57.85
+Appendix J Detailed Algorithm
+Algorithm 1 Layer-wise Adversarial Detection Measures via RT, LT, and RLT
+[…]
+[…]
+[…]
+[…]
+[…]
+6:$x$: Test input
+7:
+8:function RT($x$)
+9:  for $k=k_{RT}$ to L1L-1 do
+10:   $e_{k}\leftarrow\|z_{k}(x)-R^{(L\rightarrow k)}(z_{L}(x))\|_{2}^{2}$
+11:  end for
+12:  $\bm{e}\leftarrow(e_{1},\dots,e_{L-1})$
+13:  return $(\log(L-1)-\mathcal{H}(\sigma(\bm{e})))\cdot\log\left(\frac{1}{L-1}\sum_{k}e_{k}\right)$
+14:end function
+15:
+16:function LT($x$)
+17:  for $g=1$ to $G$ do
+18:   $\Delta z^{(g)}\leftarrow\frac{1}{L-k_{LT}+1}\sum_{i=k_{LT}}^{L}\|z_{i}(x)-z_{i}(W^{(g)}x)\|_{2}^{2}$
+19:   $\hat{y}\leftarrow\arg\max\sigma(f(x))$
+20:   $\Delta\ell^{(g)}\leftarrow\|\mathbf{o}_{\hat{y}}-\sigma(f(W^{(g)}x))\|_{2}^{2}$
+21:   $s^{(g)}\leftarrow\log\left(\mathcal{H}(\sigma(f(x)))\cdot\Delta\ell^{(g)}\right)-\log\Delta z^{(g)}$
+22:  end for
+23:  return $\frac{1}{G}\sum_{g=1}^{G}s^{(g)}$
+24:end function
+25:
+26:function RLT($x$)
+27:  $r\leftarrow$ RT($x$),  $l\leftarrow$ LT($x$)
+28:  $r_{norm}\leftarrow\Phi^{-1}(\hat{\mathcal{F}}_{RT}(r))$
+29:  $l_{norm}\leftarrow\Phi^{-1}(\hat{\mathcal{F}}_{LT}(l))$
+30:  return $r_{norm}^{2}+l_{norm}^{2}$
+31:end function